### v1

# The Impact of Feature Scaling In Machine Learning: Effects on Regression and Classification Tasks

Abstract
This research addresses the critical lack of comprehensive studies on feature scaling by systematically evaluating 12 scaling techniques - including several less common transformations - across 14 different Machine Learning algorithms and 16 datasets for classification and regression tasks. We meticulously analyzed impacts on predictive performance (using metrics such as accuracy, MAE, MSE, and R²) and computational costs (training time, inference time, and memory usage). Key findings reveal that while ensemble methods (such as Random Forest and gradient boosting models like XGBoost, CatBoost and LightGBM) demonstrate robust performance largely independent of scaling, other widely used models such as Logistic Regression, SVMs, TabNet, and MLPs show significant performance variations highly dependent on the chosen scaler. This extensive empirical analysis, with all source code, experimental results, and model parameters made publicly available to ensure complete transparency and reproducibility, offers model-specific crucial guidance to practitioners on the need for an optimal selection of feature scaling techniques.
Index Terms:
Feature Scaling, Normalization, Standardization, Data Preprocessing, Machine Learning Algorithms.
I Introduction
Machine Learning progress has been notorious in several domains of knowledge engineering, notably driven by the rise of big data [1, 2], its applications in healthcare [3, 4, 5, 6], forecasting [7, 8, 9, 10], precision agriculture [11], wireless sensor networks [12], language tasks [13, 14, 15] and many other domains [16, 17, 18, 19, 20, 21]. All different applications compose the field of Machine Learning [14], which has become a major subarea of computer science and statistics due to its crucial role in the modern world [22]. Although these methods hold immense potential for the advancement of predictive modeling, their improper application has introduced significant obstacles [23, 24, 25].
One such obstacle is the indiscriminate use of preprocessing techniques, particularly feature scaling [26]. Feature scaling is a mapping technique in a preprocessing stage by which the user tries to give all attributes the same weight [27, 28]. In some applications, this data transformation can improve the performance of Machine Learning models [29].
Consequently, applying a scaling method without a careful evaluation of its suitability for the specific problem and model may be inadvisable and could negatively impact results. This practice risks undermining the validity of the claims regarding model performance and may lead to a feedback loop of overconfidence in the results, known as overfitting [30].
Reproducibility is a critical problem in Machine Learning [31, 32, 33, 34]. It is often undermined by factors such as missing data or code, inconsistent standards, and sensitivity to training conditions [35]. Feature scaling, in particular, if not documented or applied correctly, can significantly affect model performance and hinder the replication of results. The absence of rigorous evaluation not only hampers reproducibility, but can also lead to the adoption of practices with poor generalizability across different datasets or domains.
As Machine Learning methods continue to shape research by their use in a wide range of applications, it is essential to critically assess and justify each step of the modeling pipeline, including feature scaling, to ensure robust and replicable findings [36].
The primary objective of this study is to evaluate the impact of different data scaling methods on the training process and performance metrics of various Machine Learning algorithms across multiple datasets. We employ 14 widely used Machine Learning models for tabular data, including Linear Regression, Logistic Regression, Support Vector Machines (SVM), K-Nearest Neighbors, Multilayer Perceptron, Random Forest, TabNet, Naive Bayes, Classification and Regression Trees (CART), Gradient Boosting Trees, AdaBoost, LightGBM, CatBoost, and XGBoost. These models were evaluated using 12 different data scaling techniques, in addition to a baseline without scaling, across 16 datasets covering both classification and regression tasks. The selected models represent the state of the art in tabular data analysis, offering a favorable balance between predictive performance and computational efficiency, often outperforming deep learning techniques in this context [37, 38, 39, 40].
In Section II we cover some related work in similar studies while Section III explains more about each algorithm, each feature scaling technique, the study diagram, evaluation metrics, and how the models were trained. In Section IV we represent the final results of this study and some discussion. The limitations of our study are discussed in Section V. Lastly, in Section VI we give a final conclusion of the current experiments and future works.
II Related Work
Despite its fundamental role in Machine Learning pipelines, the impact of feature scaling remains an underexplored area in the literature. Most existing studies examine only a limited number of algorithms [41] or datasets, and often provide minimal analysis of the specific effects of different scaling techniques [42] on each particular algorithm’s performance. Studies that comprehensively evaluate various scaling methods across a broad range of models and datasets, such as the approach taken in this work, are scarce. In many Machine Learning cases, preprocessing is briefly mentioned, with scaling treated as a routine step rather than a variable worthy of in-depth investigation.
For some Machine Learning models, feature scaling is extremely necessary, such as K-Nearest Neighbors [43], Neural Networks [44, 45] [46], and SVM [47] [48]. In object detection, data scaling has a crucial impact [49].
In [50], the authors compared six normalization methods in an SVM classifier to improve intrusion data, and the Min-Max Normalization showed the best performance. In [51], the authors evaluated eleven Machine Learning algorithms, including Logistic Regression, Linear Discriminant Analysis, K-Nearest Neighbors, Classification and Regression Trees, Naive Bayes, Support Vector Machine, XGBoost, Random Forest (RF), Gradient Boosting, AdaBoost, and Extra Tree, across six different data scaling methods: Normalization, Z-score Normalization, Min-Max Normalization, Max Normalization, Robust Scaler, and Quantile Transformer. However, they focused on only one dataset, the UCI - Heart Disease [52]. Despite that, their results are interesting; models based on Decision Trees showed the best performance without any scaling method, while K-Nearest Neighbors and Support Vector Machine achieved the lowest performance.
In a work on diabetes diagnosis using models such as Random Forest, Naive Bayes, K-Nearest Neighbors, Logistic Regression and Support Vector Machine [53] the researchers compared only three preprocessing scenarios: Normalization, Z-score Normalization, and no feature scaling. Their findings suggested that Random Forest, Naive Bayes, and Logistic Regression showed little sensitivity to these specific scaling approaches, and in some cases, their performance even worsened post-scaling.
Another problem is data leakage and the reproducibility of Machine Learning models. As shown in [36], many studies used preprocessing steps, such as feature scaling, on the entire dataset before splitting the data into training and test. Some studies also applied a scaling technique without knowing if that specific Machine Learning algorithm would benefit from it.
In [54] the authors demonstrate how feature scaling methods can impact the final model performance. Rather than relying on traditional scaling techniques, they propose a Generalized Logistic algorithm. This method showed particularly strong performance on datasets with a small number of samples, consistently outperforming models that used features scaled with Min-Max Normalization or Z-score Normalization.
The convergence of stochastic gradient descent is highly sensitive to the scale of input features, with studies such as [55] demonstrating that normalization is an effective method for improving convergence.
For leaf classification, a Neural Network called the Probabilistic Neural Network was used as a classifier and Min-Max normalization was applied [56]. For approval, Min-Max normalization was again used with a K-Nearest Neighbors classifier [57].
The impact of scaling and normalization techniques on NMR spectroscopic metabonomic datasets was examined in [58]. In a related study, [59] explored eight different normalization methods to enhance the biological interpretation of metabolomics data.
An in-depth study on the impact of data normalization on classification performance was presented in [41]. The authors evaluated fourteen normalization methods but employed only the K-Nearest Neighbor Classifier. Their findings indicate that normalization as a data preprocessing technique is affected by various data characteristics, such as features with differing statistical properties, the dominance of certain features, and the presence of outliers.
In an unsupervised task, [44] investigates the impact of the scaling of features on K-Means, highlighting its importance for datasets with features measured in different units. The study compares five scaling methods: Z-score Normalization, Min-Max Normalization, Percentile transformation, Maximum absolute scaling, and Robust Scaler. For cluster analysis, normalization helps prevent unwanted biases in external validation indices, such as the Jaccard, Fowlkes-Mallows, and Adjusted Rand Index, that may arise due to variations in the number of clusters or imbalances in class size distributions [60].
In the context of glaucoma detection based on a combination of texture and higher-order spectral features [61], the authors demonstrate that Z-score normalization, when paired with a Random Forest classifier, achieves superior performance compared to a Support Vector Machine.
A most recent study [62] evaluates five Machine Learning models, but focuses exclusively on classification problems and applies only two feature scaling techniques. However, most of these studies do not explain the rationale for choosing these techniques. In addition, some apply normalization to the entire dataset before splitting it into training and testing sets, leading to data leakage.
III Methodology
Our primary focus in this study is to ensure reproducibility. To that end, we use a well-known dataset for classification and regression tasks, sourced from the University of California, Irvine (UCI) Machine Learning Repository, due to its well-recognized and diverse collection of real-world datasets with standardized formats, that can easily be used for benchmarking and comparison between the different models chosen in this work.
III-A Dataset
Tables I and II provide detailed information about the datasets used in this study, including the number of features, instances, and classes. All features are numeric, represented as either int64 or float64 types. The classification tasks are either binary or multi-class.
TABLE I: Datasets used for classification: Breast Cancer Wisconsin (Diagnostic) [63], Dry Bean Dataset [64], Glass Identification [65], Heart Disease [52], Iris [66], Letter Recognition [67], MAGIC Gamma Telescope [68], Rice (Cammeo and Osmancik) [69], and Wine [70].
Dataset Instances Features Classes
Breast Cancer Wisconsin (Diagnostic) 569 30 2
Dry Bean Dataset 13611 16 7
Glass Identification 214 9 6
Heart Disease 303 13 2
Iris 150 4 3
Letter Recognition 20000 16 26
MAGIC Gamma Telescope 19020 10 2
Rice (Cammeo and Osmancik) 3810 7 2
Wine 178 13 3
TABLE II: Datasets used for regression: Air Quality [71], Abalone [72], Appliances Energy Prediction [73], Concrete Compressive Strength [74], Forest Fires [75], Real Estate Valuation [76], and Wine Quality [77].
Dataset Instances Features
Abalone 4177 8
Air Quality 9358 15
Appliances Energy Prediction 19735 28
Concrete Compressive Strength 1030 8
Forest Fires 517 12
Real Estate Valuation 414 6
Wine Quality 4898 11
III-B Train and Test set
To preserve the integrity of our analysis and avoid data leakage, we split the dataset into training and test sets prior to applying any preprocessing steps, such as feature scaling. Following standard practice in machine learning [78, 79], 70% of the data is allocated to the training set and 30% to the test set. While determining the optimal train-test split is inherently challenging in machine learning [80], our choice reflects a balance between reproducibility and the practical constraints posed by the relatively small size of some datasets. This ratio allows for both effective model training and reliable performance evaluation.
Data leakage occurs when information from the test set inadvertently influences the training process, often leading to overly optimistic performance estimates. For instance, performing oversampling or other transformations before splitting the data can introduce overlap between the training and test sets, thereby compromising their independence. By strictly separating the data prior to any preprocessing, we maintain a clear boundary between the two sets, ensuring a robust and unbiased evaluation of the model performance [36].
III-C Feature Scaling Techniques
Several feature scaling techniques were investigated. For a subset of these, we leveraged the built-in implementations available in the scikit-learn library. However, other specialized or less common scaling methods required custom implementation, which we developed as classes within our Python experimental framework.
III-C1 Min-Max Normalization (MM)
Min-Max normalization scales the data to a fixed range, typically [0, 1] [27, 81]. The transformation is given by:
$X_{\text{norm}}=\frac{X-X_{\text{min}}}{X_{\text{max}}-X_{\text{min}}},$ (1)
where $X$ is the original value, $X_{\text{min}}$ and $X_{\text{max}}$ are the minimum and maximum values in the dataset, $X_{\text{norm}}$ is the normalized value.
III-C2 Max Normalization (MA)
Max normalization scales the data by dividing each feature by its maximum absolute value [27, 82]:
$X_{\text{norm}}=\frac{X}{max(~{}|X|~{})}$ (2)
This method is advantageous when the data consists of strictly non-negative values.
III-C3 Z-score Normalization (ZSN)
Z-score normalization (also known as Standardization) transforms data to have a mean of 0 and a unit variance [27, 83]. The formula is given by:
$X_{\text{norm}}=\frac{X-\mu}{\sigma},$ (3)
where $X$ is the original feature value, $\mu$ is the mean of the feature, $\sigma$ is the standard deviation [^1] For transparency, we clarify that the population standard deviation was used throughout this study. of the feature, and $X_{\text{norm}}$ is the scaled value.
III-C4 Variable Stability Scaling (VAST)
Variable stability scaling adjusts the data based on the stability of each feature. It is particularly useful for high-dimensional datasets and can be seen as a variation of standardization that incorporates the Coefficient of Variation (CV), $\frac{\mu}{\sigma}$, as a scaling factor [59]:
$X_{\text{norm}}=\frac{(X-\mu)}{\sigma}\frac{\mu}{\sigma}$ (4)
III-C5 Pareto Scaling (PS)
Pareto scaling is a normalization technique in which each feature is centered, by subtracting the mean, and then divided by the square root of its standard deviation [84]. It’s similar to Z-score normalization, but instead of dividing by the full standard deviation, it uses its square root as the scaling factor. Pareto scaling is particularly useful when the goal is to preserve relative differences between features while reducing the impact of large variances [85, 86].
$X_{\text{norm}}=\frac{X-\mu}{\sqrt{\sigma}}$ (5)
III-C6 Mean Centered (MC)
Mean centering subtracts the mean of each feature from the data. This method is often used as a preprocessing step in Principal Component Analysis (PCA) [87].
$X_{\text{norm}}=X-\mu$ (6)
III-C7 Robust Scaler (RS)
The robust scaler uses the median and interquartile range (IQR) to scale the data:
$X_{\text{norm}}=\frac{X-X_{\text{median}}}{\text{IQR}}$
This method is robust to outliers [88].
III-C8 Quantile Transformation (QT)
Quantile transformation maps the data to a uniform or normal distribution. It is useful for non-linear data [88].
III-C9 Decimal Scaling Normalization (DS)
Decimal scaling performs normalization by adjusting the decimal point of the attribute values, thereby rescaling them to fit within the range (–1, 1), not including the endpoints [26, 27]:
$X_{\text{norm}}=\frac{X}{10^{j}}$ (7)
where $j$ is the smallest integer such that $max(~{}|X_{norm}|~{})<1$
III-C10 Tanh Transformation (TT)
A variant of tanh normalization is used, in which the Hampel estimators are replaced by the mean and standard deviation of each feature [89]:
$X_{\text{norm}}=\frac{1}{2}\left\{\tanh\left(0.01~{}\frac{X-\mu}{\sigma}\right)+1\right\}$ (8)
III-C11 Logistic Sigmoid Transformation (LS)
The logistic sigmoid-based transformation applies the sigmoid function to the data [90, 91]:
$X_{\text{norm}}=\frac{1}{1+e^{-q}},\quad\text{where}\quad q=\frac{X-\mu}{\sigma}$ (9)
III-C12 Hyperbolic Tangent Transformation (HT)
The hyperbolic tangent transformation is similar to the tanh transformation but is applied differently in certain contexts [90, 91]:
$X_{\text{norm}}=\frac{1-e^{-q}}{1+e^{-q}},\quad\text{where}\quad q=\frac{X-\mu}{\sigma}$ (10)
III-D Machine Learning Algorithms
The following models were used:
  • Logistic Regression (LR): A simple and effective statistical model for binary classification that estimates class probabilities using the logistic function [92].
  • Linear Regression: A fundamental model that fits a linear relationship between input characteristics and a continuous target variable [28, 93, 94]
  • Support Vector Machine (SVM) & Support Vector Regressor (SVR): With a linear kernel, this model finds the hyperplane that maximizes the margin between classes and for regression finds a function within a tolerance margin using the radial basis function (RBF) kernel [95, 96, 97].
  • Multilayer Perceptron (MLP): A feedforward neural network with one or more hidden layers trained using backpropagation, for classification and regression [98, 99, 100].
  • Random Forest (RF): An ensemble of decision trees built using bootstrap aggregation (bagging), improving robustness, and reducing overfitting, works for classification and regression [101, 102, 103].
  • Naive Bayes (NB): Based on Bayes’ Theorem with the assumption of conditional independence between features, only works for classification [104].
  • Classification and Regression Trees (CART): A recursive partitioning algorithm that builds binary trees for decision-making, works for classification and regression tasks [105, 106].
  • LightGBM (LGBM): A gradient boosting framework that uses histogram-based learning and leaf-wise tree growth, works for classification and regression [107].
  • AdaBoost (Ada): A boosting algorithm that iteratively focuses on misclassified examples to improve the accuracy of the model, works for classification and regression [108, 109].
  • CatBoost: A gradient boosting model with native support for categorical features, works for regression and classification [110].
  • XGBoost: An efficient implementation of gradient boosting with regularization and optimized parallel computing, works for regression and classification [111].
  • K-Nearest Neighbors (KNN): An instance-based learning algorithm that classifies samples based on the most frequent label among their nearest neighbors, also have its regression variant that predicts the target by averaging the outputs of the $k$ nearest neighbors in the feature space [112, 113, 114].
  • Attentive Interpretable Tabular Learning (TabNet): A deep learning model for tabular data that employs sequential attention to select relevant features at each decision step, works for classification and regression [115].
III-E Metrics
III-E1 Classification Metrics
  • Accuracy: one of the most widely used metrics to evaluate classification tasks. Measures the proportion of correctly predicted instances relative to the total number of predictions:
    $\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN},$ (11)
    where $TP$, $TN$, $FP$, and $FN$ represent true positives, true negatives, false positives, and false negatives, respectively.
    Despite its popularity, accuracy can be misleading when dealing with imbalanced datasets, as highlighted by [116]. However, given the characteristics of our datasets, comprising both binary and multiclass classification problems, we chose to include accuracy in our analysis, acknowledging its limitations in the presence of class imbalance.
III-E2 Regression Metrics
  • Mean Absolute Error (MAE): measures the average absolute difference between predicted and actual values, offering an intuitive sense of the magnitude of the error.
    $\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (12)
  • Mean Squared Error (MSE): calculates the average squared differences between actual and predicted values, penalizing larger errors more heavily:
    $\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (13)
  • Coefficient of Determination ($R^{2}$): indicates the proportion of variance in the dependent variable that is predictable from the independent variables. A higher value $R^{2}$ indicates a better fit of the model to the data [117].
These regression metrics are standard for evaluating continuous output predictions [118].
III-E3 Computational Metrics
To complement the evaluation of predictive performance, we also assess:
  • Memory Usage: measure memory usage during the scaling step.
  • Training Time: the time taken to train each model on a given dataset.
  • Inference Time: the time required for the trained model to make predictions on unseen data.
These metrics are essential when evaluating models for real-time or resource-constrained environments.
III-F Experiment Workflow
The objective of this subsection is to document every step of the experimentation process to ensure full transparency and reproducibility. All datasets used in this work are publicly available and can be downloaded along with their respective train/test splits. All Machine Learning algorithms were applied using their default hyperparameters, as implemented in scikit-learn or the corresponding official libraries. For algorithms that support the random_state parameter, a fixed seed was used to ensure consistent results across runs. During each training session, a configuration file is generated to record the parameters used by each model, enabling complete traceability of the experimental setup.
The experiment begins with the import and cleaning of each dataset. Categorical target variables are encoded numerically, and column names are standardized using regular expressions. After that, each dataset is partitioned into training and testing subsets, which are saved both as .csv files and as Python dictionaries. Subsequently, for every dataset and Machine Learning model combination, various scaling techniques are applied. Each model is trained on the training set and evaluated on the test set, with performance metrics computed accordingly. In addition to the results, the model configuration and metadata - including training and inference times - are stored to ensure reproducibility and to facilitate further analysis.
Algorithm 1 Experiment Workflow
1:$D$: Training Data
2:Validation metrics, training time, inference time, memory usage, and model parameters.
3:Begin
4:Import necessary libraries and load dataset $D$.
5:Perform cleaning and ETL (Extract, Transform, Load) on $D$.
6:Split $D$ into $D_{train}$ (training set) and $D_{test}$ (testing set).
7:for each available dataset do
8:     for each Machine Learning Model $M$ do
9:         Fit scaler on $D_{train}$ and apply transform to $D_{train}$ and $D_{test}$.
10:         Train Model $M$ using $D_{train}$.
11:         Evaluate $M$ on $D_{test}$.
12:         Save validation results, including accuracy, training time, inference time, memory usage, and model parameters.
13:     end for
14:end for
15:End
III-G Python Script Descriptions
The following scripts were developed to automate and manage the experimental pipeline:
  • import_dataset.py: Imports datasets from the UCI repository and maps the appropriate target variable for each dataset.
  • etl_cleaning.py: Convert categorical and numerical variables as needed and cleans column names using regular expressions.
  • train_test_split.py: Splits the datasets into training and testing sets and saves them in both .csv and dictionary formats.
  • train_results.py: Train each Machine Learning model on every dataset using different scaling techniques. Calculate validation metrics and save both performance results and model configuration.
  • main.py: Serves as the main execution script that orchestrates all stages of the experiment.
Figure 1: Experimental design
III-H Source Code of this Experiments
For complete transparency and reproducibility, the source code, all experimental results, and detailed model parameters are publicly available in GitHub
The experiments were conducted on a system running a 64-bit Linux distribution, equipped with an AMD Ryzen™ 9 7900 processor (capable of boosting to 5.4 GHz) and 64 GB of RAM.
IV Experiments
This section presents the empirical results of our extensive experiments. We selected five models and three scaling techniques, along with a baseline without scaling, as this subset already allows us to draw meaningful conclusions and observe differences between models. The complete results are available in the Appendix A.
IV-A Impact on Validation Metrics
TABLE III: Accuracy results by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVM under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Breast Cancer Wisconsin Diagnostic KNN 0.9591 0.9766 0.9591 0.9649
LGBM 0.9474 0.9474 0.9591 0.9591
MLP 0.9649 0.9766 0.9766 0.9708
RF 0.9708 0.9708 0.9708 0.9708
SVM 0.9240 0.9825 0.9766 0.9825
Dry Bean KNN 0.7113 0.9141 0.9216 0.9190
LGBM 0.9275 0.9275 0.9263 0.9275
MLP 0.2980 0.9101 0.9327 0.9314
RF 0.9238 0.9226 0.9226 0.9226
SVM 0.5803 0.9109 0.9263 0.9268
Glass Identification KNN 0.5846 0.6615 0.6308 0.6308
LGBM 0.8154 0.8154 0.8000 0.8308
MLP 0.7077 0.7231 0.6923 0.6769
RF 0.7538 0.7538 0.7692 0.7846
SVM 0.6769 0.5077 0.6615 0.6462
Heart Disease KNN 0.4945 0.5495 0.5714 0.5275
LGBM 0.5275 0.5275 0.5275 0.5385
MLP 0.3516 0.5495 0.5055 0.5385
RF 0.5604 0.5604 0.5604 0.5604
SVM 0.5055 0.5934 0.6154 0.5934
Iris KNN 1.0000 1.0000 1.0000 0.9556
LGBM 1.0000 1.0000 1.0000 1.0000
MLP 1.0000 1.0000 1.0000 1.0000
RF 1.0000 1.0000 1.0000 1.0000
SVM 1.0000 1.0000 0.9778 0.9778
Letter Recognition KNN 0.9493 0.9480 0.9405 0.9158
LGBM 0.9640 0.9640 0.9637 0.9637
MLP 0.9367 0.9280 0.9502 0.9545
RF 0.9577 0.9577 0.9570 0.9580
SVM 0.8135 0.8208 0.8488 0.8488
Magic Gamma Telescope KNN 0.8098 0.8254 0.8340 0.8340
LGBM 0.8803 0.8803 0.8792 0.8810
MLP 0.8170 0.8677 0.8717 0.8777
RF 0.8808 0.8808 0.8808 0.8808
SVM 0.2976 0.5158 0.4341 0.3212
Rice Cammeo And Osmancik KNN 0.8775 0.9204 0.9143 0.9064
LGBM 0.9213 0.9213 0.9178 0.9160
MLP 0.5468 0.9335 0.9318 0.9309
RF 0.9265 0.9265 0.9265 0.9265
SVM 0.9248 0.9309 0.9309 0.9274
Wine KNN 0.7407 0.9444 0.9630 0.9444
LGBM 0.9815 0.9815 0.9815 0.9815
MLP 0.9815 1.0000 0.9815 0.9815
RF 1.0000 1.0000 1.0000 1.0000
SVM 0.5926 1.0000 0.9815 0.9815
TABLE IV: $R^{2}$ score by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVR under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Abalone KNN 0.5164 0.4955 0.4662 0.4552
LGBM 0.5260 0.5260 0.5256 0.5190
MLP 0.5245 0.5265 0.5578 0.5632
RF 0.5244 0.5249 0.5234 0.5241
SVR 0.5293 0.5257 0.5421 0.5398
Air Quality KNN 0.9995 0.9993 0.9994 0.9986
LGBM 0.9999 0.9999 0.9999 0.9999
MLP 0.9985 1.0000 1.0000 0.9999
RF 1.0000 1.0000 1.0000 1.0000
SVR 0.9966 0.9619 0.9269 0.9188
Appliances Energy Prediction KNN 0.1681 0.2049 0.3279 0.2929
LGBM 0.4318 0.4318 0.4334 0.4192
MLP 0.1598 0.1581 0.3144 0.2970
RF 0.5122 0.5120 0.5122 0.5125
SVR -0.1056 -0.0275 0.0154 -0.0096
Concrete Compressive Strength KNN 0.6770 0.6631 0.6714 0.7446
LGBM 0.9229 0.9229 0.9217 0.9226
MLP 0.8030 0.7468 0.8725 0.8662
RF 0.8896 0.8895 0.8891 0.8894
SVR 0.2259 0.5394 0.6093 0.6987
Forest Fires KNN -0.0115 -0.0470 -0.0447 -0.0345
LGBM -0.0246 -0.0246 -0.0188 -0.0135
MLP -0.0067 0.0057 0.0013 0.0121
RF -0.1060 -0.1101 -0.1093 -0.1058
SVR -0.0257 -0.0246 -0.0244 -0.0245
Real Estate Valuation KNN 0.6232 0.6232 0.6153 0.6348
LGBM 0.7001 0.7001 0.7120 0.7075
MLP 0.6199 0.5608 0.6436 0.6894
RF 0.7444 0.7449 0.7444 0.7444
SVR 0.4897 0.5327 0.5788 0.5872
Wine Quality KNN 0.1221 0.3283 0.3475 0.3234
LGBM 0.4557 0.4557 0.4578 0.4602
MLP 0.2500 0.3243 0.3872 0.3879
RF 0.4985 0.4991 0.4982 0.4992
SVR 0.1573 0.3185 0.3842 0.3799
TABLE V: MSE by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVR under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Abalone KNN 4.9107 5.1229 5.4200 5.5324
LGBM 4.8137 4.8137 4.8170 4.8846
MLP 4.8285 4.8087 4.4900 4.4356
RF 4.8295 4.8244 4.8392 4.8328
SVR 4.7801 4.8158 4.6502 4.6734
Air Quality KNN 0.9109 1.2492 1.0231 2.3434
LGBM 0.1129 0.1129 0.1298 0.1291
MLP 2.6090 0.0278 0.0110 0.1181
RF 0.0131 0.0131 0.0131 0.0131
SVR 5.7798 65.6306 125.8177 139.7765
Appliances Energy Prediction KNN 8570.5671 8192.0770 6924.2331 7285.3653
LGBM 5854.3508 5854.3508 5837.8269 5983.4385
MLP 8656.1894 8673.7742 7063.7610 7243.1268
RF 5025.6175 5028.0530 5025.4432 5022.6680
SVR 11390.9162 10585.9003 10144.3973 10401.7762
Concrete Compressive Strength KNN 87.4024 91.1487 88.9093 69.1034
LGBM 20.8520 20.8520 21.1791 20.9335
MLP 53.3124 68.5215 34.5083 36.1983
RF 29.8643 29.8945 30.0114 29.9231
SVR 209.4406 124.6258 105.7187 81.5215
Forest Fires KNN 8049.5193 8331.5751 8314.0348 8232.2422
LGBM 8153.9240 8153.9240 8107.4363 8065.4234
MLP 8010.9478 7912.5614 7947.6005 7861.6344
RF 8801.7409 8833.7029 8827.5833 8799.6126
SVR 8162.5768 8154.0176 8152.2006 8152.5913
Real Estate Valuation KNN 63.0027 63.0182 64.3252 61.0770
LGBM 50.1547 50.1547 48.1590 48.9211
MLP 63.5585 73.4388 59.6005 51.9328
RF 42.7505 42.6657 42.7479 42.7405
SVR 85.3424 78.1453 70.4354 69.0369
Wine Quality KNN 0.6406 0.4901 0.4761 0.4937
LGBM 0.3971 0.3971 0.3956 0.3939
MLP 0.5472 0.4930 0.4471 0.4466
RF 0.3659 0.3655 0.3661 0.3654
SVR 0.6149 0.4973 0.4493 0.4525
TABLE VI: MAE by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVR under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Abalone KNN 1.5673 1.6102 1.6555 1.6654
LGBM 1.5476 1.5476 1.5500 1.5632
MLP 1.6159 1.5748 1.5318 1.4958
RF 1.5590 1.5584 1.5617 1.5619
SVR 1.5048 1.5111 1.4964 1.4990
Air Quality KNN 0.5506 0.6755 0.5862 0.8379
LGBM 0.0677 0.0677 0.0728 0.0731
MLP 1.2505 0.0909 0.0698 0.1910
RF 0.0167 0.0167 0.0166 0.0167
SVR 0.8970 1.4802 1.9293 5.4424
Appliances Energy Prediction KNN 47.7696 45.5977 39.8189 41.5828
LGBM 39.0633 39.0633 39.0969 39.3198
MLP 53.8080 54.1215 47.7318 47.7236
RF 34.2701 34.2923 34.2876 34.2838
SVR 48.9182 45.5442 43.4164 45.1344
Concrete Compressive Strength KNN 7.2301 7.0405 7.3319 6.4237
LGBM 3.0480 3.0480 3.0473 3.0197
MLP 5.9685 6.4045 4.4758 4.5392
RF 3.7512 3.7503 3.7608 3.7550
SVR 11.6674 8.9978 8.1314 7.0217
Forest Fires KNN 21.5065 20.7837 21.0229 19.7818
LGBM 24.2821 24.2821 24.1922 23.9371
MLP 21.2084 20.6192 24.5951 23.9833
RF 24.2812 24.3738 24.3496 24.3443
SVR 14.9474 14.9755 14.9655 14.9590
Real Estate Valuation KNN 5.4626 5.4435 5.6979 5.5819
LGBM 4.8106 4.8106 4.7284 4.8254
MLP 5.3601 6.1126 5.4400 5.0138
RF 4.3971 4.3951 4.4002 4.3916
SVR 6.8824 6.2845 5.9805 5.8563
Wine Quality KNN 0.6243 0.5231 0.5259 0.5349
LGBM 0.4832 0.4832 0.4829 0.4833
MLP 0.5692 0.5499 0.5187 0.5196
RF 0.4366 0.4368 0.4370 0.4362
SVR 0.6076 0.5453 0.5111 0.5142
As anticipated, a key finding from our classification experiments is the differential impact of feature scaling on model performance. Table III shows the accuracy results. Ensemble methods, including Random Forest and the gradient boosting family (LightGBM, CatBoost, XGBoost), proved robust, delivering consistently high validation metrics irrespective of the preprocessing strategy. This inherent robustness offers a significant practical advantage, particularly in resource-constrained environments, since omitting the scaling step eliminates the associated memory and computational overhead. The Naive Bayes model showed similar scaling resistance, although its overall accuracy was not competitive with these top-tier ensembles. In stark contrast, the performance of Logistic Regression (LR), Support Vector Machines (SVM), K-Nearest Neighbor (KNN), TabNet and Multi-Layer Perceptrons (MLP) was highly dependent on the choice of scaler, revealing their pronounced sensitivity to data preprocessing through significant fluctuations in performance.
A similar pattern of scaling sensitivity was observed in regression tasks when employing the regression counterparts of these classification models, as shows Tables IV, V and VI. This suggests that the underlying mathematical principles governing these model families lead to consistent behavior regarding data scaling, whether applied to classification or regression problems.
In general, our findings affirm the superior performance of the ensemble methods, Random Forest, LightGBM, CatBoost, and XGBoost, which is consistent with their established reputation as state-of-the-art models for tabular data. Their ability to achieve high accuracy regardless of the scaling technique applied explains why feature scaling is often considered an optional preprocessing step for these particular models in many Machine Learning projects. This practical consideration, combined with their predictive power, underscores their utility in a wide range of applications.
IV-B Train and Inference time Impact
TABLE VII: Time to Train (s) in classification models by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVM under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Breast Cancer Wisconsin Diagnostic KNN 0.0002 0.0002 0.0002 0.0002
LGBM 0.0213 0.0219 0.0275 0.0256
MLP 0.1708 0.2929 0.1683 0.1824
RF 0.0706 0.0707 0.0709 0.0706
SVM 0.0016 0.0007 0.0009 0.0009
Dry Bean KNN 0.0008 0.0008 0.0008 0.0008
LGBM 0.2732 0.2759 0.2988 0.3025
MLP 0.3396 3.2012 2.4228 3.8822
RF 1.8703 1.8860 1.8872 1.8734
SVM 0.1096 0.1932 0.1560 0.1470
Glass Identification KNN 0.0002 0.0002 0.0002 0.0004
LGBM 0.0422 0.0427 0.0345 0.0410
MLP 0.2307 0.2416 0.2376 0.2372
RF 0.0453 0.0456 0.0453 0.0453
SVM 0.0010 0.0007 0.0008 0.0009
Heart Disease KNN 0.0003 0.0003 0.0003 0.0003
LGBM 0.0466 0.0450 0.0412 0.0439
MLP 0.0198 0.1105 0.3423 0.3942
RF 0.0449 0.0452 0.0451 0.0451
SVM 0.0029 0.0009 0.0017 0.0012
Iris KNN 0.0004 0.0002 0.0002 0.0003
LGBM 0.0116 0.0119 0.0130 0.0118
MLP 0.1074 0.1444 0.0947 0.1164
RF 0.0382 0.0387 0.0386 0.0385
SVM 0.0003 0.0004 0.0003 0.0003
Letter Recognition KNN 0.0010 0.0009 0.0009 0.0010
LGBM 1.1319 1.0970 1.1101 1.1380
MLP 11.1298 22.0963 11.5231 13.5357
RF 0.8526 0.8582 0.8632 0.8535
SVM 0.8484 0.7841 0.7588 0.7936
Magic Gamma Telescope KNN 0.0065 0.0065 0.0065 0.0071
LGBM 0.0469 0.0460 0.0461 0.0468
MLP 0.5592 4.8919 4.0988 3.4037
RF 2.6338 2.6361 2.6471 2.6319
SVM 0.0697 0.2742 0.2290 0.2240
Rice Cammeo And Osmancik KNN 0.0009 0.0010 0.0010 0.0010
LGBM 0.0294 0.0291 0.0309 0.0315
MLP 0.0464 0.6241 0.1954 0.2864
RF 0.2023 0.2032 0.2026 0.2034
SVM 0.0093 0.0223 0.0201 0.0199
Wine KNN 0.0003 0.0003 0.0002 0.0003
LGBM 0.0164 0.0151 0.0167 0.0154
MLP 0.2105 0.1658 0.0603 0.0696
RF 0.0406 0.0408 0.0411 0.0406
SVM 0.0008 0.0004 0.0005 0.0005
TABLE VIII: Time to Inference (s) in classification models by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVM under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Breast Cancer Wisconsin Diagnostic KNN 0.0025 0.0025 0.0025 0.0030
LGBM 0.0004 0.0004 0.0007 0.0007
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0013 0.0013 0.0013 0.0013
SVM 0.0002 0.0002 0.0001 0.0001
Dry Bean KNN 0.0529 0.0536 0.0536 0.0594
LGBM 0.0101 0.0099 0.0099 0.0100
MLP 0.0011 0.0012 0.0012 0.0012
RF 0.0180 0.0181 0.0184 0.0179
SVM 0.0255 0.1723 0.0892 0.0907
Glass Identification KNN 0.0012 0.0012 0.0012 0.0014
LGBM 0.0005 0.0006 0.0006 0.0006
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0013 0.0013 0.0013 0.0013
SVM 0.0002 0.0002 0.0002 0.0002
Heart Disease KNN 0.0016 0.0016 0.0016 0.0020
LGBM 0.0007 0.0006 0.0006 0.0006
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0014 0.0014 0.0014 0.0015
SVM 0.0003 0.0002 0.0002 0.0002
Iris KNN 0.0012 0.0010 0.0009 0.0012
LGBM 0.0004 0.0005 0.0004 0.0004
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0011 0.0011 0.0011 0.0011
SVM 0.0001 0.0001 0.0001 0.0001
Letter Recognition KNN 0.8666 0.0819 0.0821 0.0870
LGBM 0.0611 0.0606 0.0610 0.0601
MLP 0.0022 0.0022 0.0021 0.0021
RF 0.0471 0.0473 0.0472 0.0479
SVM 0.8372 1.3598 0.8540 0.8722
Magic Gamma Telescope KNN 0.1088 0.1550 0.1659 0.1635
LGBM 0.0024 0.0023 0.0024 0.0023
MLP 0.0012 0.0012 0.0012 0.0012
RF 0.0347 0.0349 0.0349 0.0341
SVM 0.0214 0.1027 0.0838 0.0856
Rice Cammeo And Osmancik KNN 0.0139 0.0148 0.0149 0.0152
LGBM 0.0007 0.0007 0.0007 0.0008
MLP 0.0003 0.0003 0.0002 0.0003
RF 0.0045 0.0044 0.0045 0.0044
SVM 0.0012 0.0083 0.0055 0.0056
Wine KNN 0.0011 0.0011 0.0011 0.0011
LGBM 0.0004 0.0004 0.0004 0.0004
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0011 0.0011 0.0011 0.0011
SVM 0.0001 0.0001 0.0001 0.0001
TABLE IX: Time to Train (s) in regression models by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVM under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Abalone KNN 0.0010 0.0010 0.0010 0.0010
LGBM 0.0286 0.0288 0.0301 0.0307
MLP 0.8414 1.1404 0.8932 1.2343
RF 0.5945 0.6035 0.6003 0.5977
SVR 0.1382 0.1383 0.1392 0.1440
Air Quality KNN 0.0032 0.0032 0.0031 0.0031
LGBM 0.0391 0.0395 0.0392 0.0396
MLP 0.5701 3.3321 2.1864 1.8880
RF 1.8482 1.8664 1.8547 1.8659
SVR 0.6315 0.5742 0.4535 0.7643
Appliances Energy Prediction KNN 0.0004 0.0004 0.0004 0.0004
LGBM 0.0522 0.0532 0.0552 0.0545
MLP 3.3408 12.6484 17.8381 16.1881
RF 18.2580 18.3987 18.3023 18.2253
SVR 3.6868 3.6063 3.6767 3.7207
Concrete Compressive Strength KNN 0.0004 0.0004 0.0004 0.0004
LGBM 0.0221 0.0232 0.0231 0.0237
MLP 0.1335 0.7281 0.8502 0.8323
RF 0.1397 0.1396 0.1390 0.1438
SVR 0.0086 0.0093 0.0087 0.0091
Forest Fires KNN 0.0003 0.0003 0.0003 0.0003
LGBM 0.0115 0.0121 0.0119 0.0118
MLP 0.2457 0.3816 0.4316 0.4330
RF 0.0997 0.1006 0.1002 0.1063
SVR 0.0025 0.0028 0.0027 0.0025
Real Estate Valuation KNN 0.0002 0.0002 0.0002 0.0002
LGBM 0.0090 0.0094 0.0094 0.0098
MLP 0.1022 0.3386 0.3599 0.3778
RF 0.0664 0.0666 0.0663 0.0689
SVR 0.0019 0.0018 0.0023 0.0017
Wine Quality KNN 0.0022 0.0022 0.0023 0.0022
LGBM 0.0314 0.0313 0.0322 0.0334
MLP 0.4713 1.3455 2.3311 1.6345
RF 1.2844 1.2891 1.2968 1.3414
SVR 0.3166 0.3181 0.3239 0.3213
TABLE X: Time to Inference (s) in regression models by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVM under the None, MA, ZSN, and RS scaling methods. Full results are provided in the Appendix A.
Dataset Model NO MA ZSN RS
Abalone KNN 0.0034 0.0048 0.0042 0.0041
LGBM 0.0006 0.0007 0.0007 0.0007
MLP 0.0003 0.0003 0.0003 0.0003
RF 0.0110 0.0111 0.0110 0.0109
SVR 0.0652 0.0652 0.0656 0.0666
Air Quality KNN 0.0171 0.0317 0.0338 0.0381
LGBM 0.0010 0.0009 0.0009 0.0009
MLP 0.0005 0.0006 0.0006 0.0006
RF 0.0169 0.0171 0.0170 0.0171
SVR 0.2986 0.2676 0.1981 0.3599
Appliances Energy Prediction KNN 0.0222 0.0215 0.0225 0.0216
LGBM 0.0017 0.0017 0.0017 0.0017
MLP 0.0012 0.0012 0.0012 0.0012
RF 0.0694 0.0697 0.0700 0.0692
SVR 1.8656 1.8341 1.7995 1.8159
Concrete Compressive Strength KNN 0.0009 0.0010 0.0010 0.0010
LGBM 0.0004 0.0004 0.0004 0.0004
MLP 0.0001 0.0002 0.0001 0.0001
RF 0.0033 0.0034 0.0032 0.0034
SVR 0.0044 0.0043 0.0041 0.0041
Forest Fires KNN 0.0003 0.0006 0.0006 0.0005
LGBM 0.0003 0.0004 0.0003 0.0003
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0018 0.0018 0.0018 0.0018
SVR 0.0011 0.0012 0.0011 0.0013
Real Estate Valuation KNN 0.0003 0.0003 0.0003 0.0003
LGBM 0.0003 0.0003 0.0003 0.0003
MLP 0.0001 0.0001 0.0001 0.0001
RF 0.0017 0.0017 0.0017 0.0017
SVR 0.0007 0.0009 0.0007 0.0007
Wine Quality KNN 0.0048 0.0257 0.0369 0.0329
LGBM 0.0008 0.0008 0.0008 0.0009
MLP 0.0004 0.0004 0.0004 0.0005
RF 0.0143 0.0147 0.0145 0.0151
SVR 0.1501 0.1527 0.1419 0.1422
The application of different scaling techniques had a variable impact on inference times in all models. For training time, the effects of scaling largely mirrored those seen in the validation accuracy results, as shown in Tables VII and IX. Certain models, notably tree-based ensembles, did not derive a significant speed benefit from feature scaling, while others were more sensitive. The preprocessing step for inference generally introduced a small, non-uniform computational cost, which could become more pronounced for large datasets or in applications requiring real-time inference.
IV-C Results of Memory Usage (kB)
The analysis showed a clear distinction in memory consumption between scaling techniques. As expected, applying no scaling resulted in zero additional memory consumption, only the consumption to load the data. Among the actual scaling methods, the RobustScaler, StandardScaler,Tanh Transformer, and Hyperbolic Tangent were found to be the most memory intensive. In contrast, the MaxAbsScaler, MinMaxScaler, and Decimal Scaler consistently registered the lowest memory usage as shows in Table XI.
TABLE XI: Memory Usage (kB) per Dataset and Scaling Method
Dataset NO MA ZSN RS
Breast Cancer
Wisconsin
Diagnostic
0.1875 175.7594 176.1266 384.9979
Dry Bean 1704.2750 2448.3812 2599.3156 2388.1666
Glass
Identification
0.1875 23.1984 26.1799 51.8070
Heart Disease 33.6609 67.2016 71.7602 122.0010
Iris 0.1875 8.7828 10.5523 20.2666
Letter
Recognition
0.1875 3566.3000 3787.1906 2568.4697
Magic Gamma
Telescope
0.1875 1552.0641 1552.2750 1552.7039
Rice Cammeo
And Osmancik
211.2688 358.2641 378.0189 297.6197
Wine 21.0234 40.3781 43.8195 73.9416
Abalone 0.1875 294.4004 294.5879 294.9893
Air Quality 880.1270 1294.5430 1373.1387 1233.9355
Appliances
Energy
Prediction
4164.2959 5894.4902 6261.5137 5832.6855
Concrete
Compressive
Strength
65.6523 137.7617 144.9668 94.3340
Forest Fires 43.2832 87.1992 92.4219 157.8330
Real Estate
Valuation
22.2676 43.1992 46.3398 79.0088
Wine Quality 0.1875 624.3691 624.5879 625.0762
V Limitation
While this study provides a broad empirical analysis of feature scaling across various models and datasets, certain limitations should be acknowledged, which also open avenues for future research.
  • Hyperparameter Optimization: The Machine Learning models in this study were evaluated using their default hyperparameters, as stated in the methodology. A comprehensive hyperparameter tuning process for each model-scaler-dataset combination was beyond the current scope but could potentially reveal different optimal pairings or further enhance model performance.
  • Scope and Diversity of datasets: Although 16 datasets were used for both classification and regression tasks, the findings could be further enhanced by incorporating an even wider array of datasets, particularly those with very high dimensionality, different types of underlying data distributions, or from more specialized domains.
  • Evaluation Metrics for Classification: The primary metric for classification tasks was accuracy. Although acknowledged as potentially misleading for imbalanced datasets, future work could incorporate a broader suite of metrics, such as F1 score, precision-recall AUC, or balanced precision, to provide a more nuanced understanding of performance, especially on datasets with skewed class distributions.
  • Dataset Size and Synthetic Data: For some of the smaller datasets utilized, the exploration of techniques like synthetic data generation or data augmentation was not performed. Such methods could potentially improve the robustness and performance of certain models, particularly data-hungry algorithms, and would be an interesting direction for further study.
  • Focus on Default Algorithm Implementations: The study relied on standard implementations of algorithms mainly from well-known libraries. Investigating variations or more recent advancements within these algorithm families could offer additional insight.
These limitations are common in empirical studies of this nature and primarily highlight areas where this already extensive work could be expanded in the future.
VI Conclusion
This comprehensive empirical study investigated the impact of 12 feature scaling techniques on 13 Machine Learning algorithms in 16 classification and regression datasets. Key findings reaffirmed the robustness of ensemble methods (e.g., Random Forest, gradient boosting family), which, along with Naive Bayes, largely maintained high performance irrespective of scaling. This offers efficiency gains by potentially avoiding preprocessing overhead. In stark contrast, models such as logistic regression, SVMs, MLPs, K-Nearest Neighbor and TabNet demonstrated high sensitivity, with their performance critically dependent on scaler choice - a pattern consistent across both task types. Computational analysis also indicated that scaling choices can influence training/inference times and memory usage, with certain scalers being notably more resource-intensive.
This study contributes with one of the first systematic evaluations of such an extensive array of models, some less common models like TabNet, and scaling techniques—including transformations, e.g Tanh (TT) and Hyperbolic Tangent (HT), which are less commonly benchmarked as general-purpose scalers, all within a unified Python framework. This is particularly relevant given that feature scaling is often applied in the literature without clear rationale, sometimes incorrectly before data splitting, which can lead to data leakage, or without verifying algorithm-specific benefits. By providing broad empirical evidence, our work offers clear guidance on how to mitigate these common issues, promoting informed scaling selection and more rigorous experimental design.
Future research could extend these insights by exploring extensive hyperparameter optimization, incorporating more diverse datasets, and utilizing a broader suite of evaluation metrics. Nevertheless, this study significantly contributes to a deeper, practical understanding of feature scaling’s role in Machine Learning.
Appendix A Tables Results
TABLE XII: Accuracy by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
Breast Cancer Wisconsin Diagnostic Ada 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766
CART 0.9415 0.9415 0.9415 0.9415 0.9415 0.9415 0.9415 0.9415 0.9474 0.9415 0.9415 0.9474 0.9474
CatBoost 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766 0.9766
KNN 0.9591 0.9649 0.9766 0.9591 0.9766 0.9591 0.9591 0.9649 0.9649 0.9474 0.9591 0.9591 0.9591
LGBM 0.9474 0.9474 0.9474 0.9591 0.9591 0.9591 0.9591 0.9591 0.9474 0.9474 0.9415 0.9415 0.9591
LR 0.9708 0.9649 0.9532 0.9825 0.9766 0.9883 0.9708 0.9883 0.9942 0.9123 0.6316 0.9883 0.9942
MLP 0.9649 0.9766 0.9766 0.9766 0.9883 0.9708 0.9006 0.9708 0.9766 0.9825 0.6316 0.9825 0.9883
NB 0.9415 0.9357 0.9357 0.9357 0.9357 0.9357 0.9415 0.9357 0.9532 0.9357 0.9357 0.9532 0.9532
RF 0.9708 0.9708 0.9708 0.9708 0.9708 0.9708 0.9708 0.9708 0.9708 0.9708 0.9708 0.9708 0.9708
SVM 0.9240 0.9883 0.9825 0.9766 0.9766 0.9766 0.9591 0.9825 0.9766 0.9532 0.6316 0.9883 0.9883
TabNet 0.5673 0.7310 0.6608 0.6608 0.7544 0.5906 0.7018 0.6433 0.6140 0.6023 0.3684 0.5263 0.7076
XGBoost 0.9649 0.9649 0.9649 0.9649 0.9649 0.9649 0.9649 0.9649 0.9649 0.9649 0.9649 0.9649 0.9649
Dry Bean Ada 0.6467 0.6467 0.6467 0.6467 0.6467 0.6467 0.6467 0.6467 0.6467 0.6467 0.6467 0.6467 0.6467
CART 0.8881 0.8881 0.8881 0.8881 0.8881 0.8881 0.8881 0.8881 0.8879 0.8881 0.8881 0.8881 0.8881
CatBoost 0.9305 0.9297 0.9300 0.9297 0.9297 0.9297 0.9297 0.9297 0.9297 0.9305 0.9319 0.9297 0.9297
KNN 0.7113 0.9197 0.9141 0.9216 0.8984 0.7507 0.7113 0.9190 0.9192 0.9158 0.9216 0.9229 0.9229
LGBM 0.9275 0.9275 0.9275 0.9263 0.9263 0.9263 0.9263 0.9275 0.9275 0.9275 0.9275 0.9275 0.9263
LR 0.7054 0.9185 0.9070 0.9229 0.9089 0.8805 0.6876 0.9234 0.9175 0.9070 0.3127 0.9246 0.9251
MLP 0.2980 0.9224 0.9101 0.9327 0.9119 0.9199 0.5992 0.9314 0.9319 0.9087 0.8834 0.9297 0.9334
NB 0.7627 0.8999 0.8999 0.8999 0.9003 0.8999 0.7627 0.8999 0.8905 0.8999 0.8999 0.8979 0.8979
RF 0.9238 0.9226 0.9226 0.9226 0.9226 0.9226 0.9238 0.9226 0.9231 0.9226 0.9226 0.9231 0.9231
SVM 0.5803 0.9246 0.9109 0.9263 0.5553 0.6060 0.1379 0.9268 0.9273 0.9111 0.7253 0.9253 0.9263
TabNet 0.8763 0.9285 0.9170 0.9273 0.9209 0.9224 0.9038 0.9243 0.9155 0.9197 0.5228 0.9246 0.9258
XGBoost 0.9261 0.9273 0.9256 0.9273 0.9273 0.9273 0.9273 0.9273 0.9265 0.9261 0.9248 0.9265 0.9273
Glass Identification Ada 0.5077 0.5077 0.5077 0.5077 0.5077 0.5077 0.5077 0.5077 0.5077 0.5077 0.5077 0.5077 0.5077
CART 0.6462 0.6462 0.6462 0.6462 0.6462 0.6462 0.6462 0.6462 0.6462 0.6462 0.6462 0.6462 0.6462
CatBoost 0.7846 0.7846 0.7846 0.7846 0.7846 0.7846 0.7846 0.7846 0.7846 0.7846 0.7846 0.7846 0.7846
KNN 0.5846 0.5846 0.6615 0.6308 0.6308 0.5385 0.5846 0.6308 0.6615 0.6462 0.6308 0.6923 0.6923
LGBM 0.8154 0.8000 0.8154 0.8000 0.8000 0.8000 0.8000 0.8308 0.8000 0.8154 0.8154 0.8154 0.8000
LR 0.6615 0.5231 0.5077 0.6769 0.7077 0.5385 0.6769 0.6615 0.6615 0.4308 0.3538 0.6000 0.6462
MLP 0.7077 0.7231 0.7231 0.6923 0.6615 0.7385 0.6615 0.6769 0.7538 0.6462 0.3538 0.7538 0.6462
NB 0.3077 0.3077 0.3077 0.3077 0.3077 0.4923 0.3077 0.3077 0.2615 0.3077 0.3077 0.3231 0.3231
RF 0.7538 0.7538 0.7538 0.7692 0.7538 0.7692 0.7538 0.7846 0.8000 0.7846 0.7692 0.7692 0.7692
SVM 0.6769 0.6308 0.5077 0.6615 0.6615 0.3231 0.6769 0.6462 0.6769 0.4923 0.3538 0.6615 0.6462
TabNet 0.1692 0.3231 0.0923 0.2154 0.1846 0.1538 0.3231 0.1692 0.2923 0.2923 0.2923 0.2923 0.2615
XGBoost 0.7692 0.7692 0.7692 0.7692 0.7692 0.7692 0.7692 0.7692 0.7692 0.7692 0.7692 0.7692 0.7692
Heart Disease Ada 0.5385 0.5385 0.5385 0.5385 0.5385 0.5385 0.5385 0.5385 0.5495 0.5385 0.5385 0.5385 0.5385
CART 0.4176 0.4176 0.4176 0.4176 0.4176 0.4176 0.4176 0.4176 0.4176 0.4176 0.4176 0.4176 0.4176
CatBoost 0.5714 0.5714 0.5714 0.5714 0.5714 0.5714 0.5714 0.5714 0.5714 0.5714 0.5714 0.5714 0.5714
KNN 0.4945 0.5275 0.5495 0.5714 0.5165 0.4835 0.4945 0.5275 0.5714 0.5824 0.5714 0.5495 0.5495
LGBM 0.5275 0.5275 0.5275 0.5275 0.5275 0.5275 0.5275 0.5385 0.5275 0.5275 0.5275 0.5275 0.5275
LR 0.5934 0.5824 0.5934 0.5385 0.5495 0.5604 0.5495 0.5714 0.5714 0.5604 0.5275 0.5934 0.5824
MLP 0.3516 0.5824 0.5495 0.5055 0.5055 0.4945 0.5604 0.5385 0.5275 0.5604 0.5275 0.5824 0.5385
NB 0.4396 0.3187 0.3187 0.2967 0.3846 0.3846 0.4396 0.3407 0.3187 0.3187 0.2967 0.3297 0.3297
RF 0.5604 0.5604 0.5604 0.5604 0.5604 0.5604 0.5604 0.5604 0.5385 0.5604 0.5604 0.5604 0.5604
SVM 0.5055 0.6154 0.5934 0.6154 0.5275 0.5385 0.4945 0.5934 0.5934 0.5714 0.5275 0.5934 0.5934
TabNet 0.3187 0.1099 0.0989 0.1099 0.2418 0.2198 0.1538 0.1758 0.1209 0.1099 0.1099 0.1209 0.1538
XGBoost 0.5165 0.5165 0.5165 0.5165 0.5165 0.5165 0.5165 0.5165 0.5165 0.5165 0.5165 0.5165 0.5165
Iris Ada 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
CART 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 0.9556 1.0000 1.0000 0.9556 0.9556
CatBoost 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
KNN 1.0000 1.0000 1.0000 1.0000 1.0000 0.9111 1.0000 0.9556 1.0000 1.0000 1.0000 1.0000 1.0000
LGBM 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
LR 1.0000 0.9111 0.9333 1.0000 1.0000 1.0000 1.0000 1.0000 0.9111 0.8222 0.4000 0.8667 0.9333
MLP 1.0000 0.9778 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 0.2889 1.0000 1.0000
NB 0.9778 0.9778 0.9778 0.9778 0.9778 0.9778 0.9778 0.9778 0.9778 0.9778 0.9778 0.9778 0.9778
RF 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
SVM 1.0000 1.0000 1.0000 0.9778 1.0000 0.9778 1.0000 0.9778 1.0000 0.9778 0.4222 0.9556 1.0000
TabNet 0.4222 0.4667 0.3778 0.2667 0.3111 0.4000 0.2667 0.2222 0.4000 0.4000 0.2889 0.6222 0.3111
XGBoost 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
Letter Recognition Ada 0.2378 0.2378 0.2378 0.2378 0.2378 0.2378 0.2378 0.2378 0.2378 0.2378 0.2378 0.2378 0.2378
CART 0.8760 0.8745 0.8745 0.8748 0.8748 0.8750 0.8748 0.8765 0.8747 0.8752 0.8753 0.8755 0.8755
CatBoost 0.9640 0.9640 0.9640 0.9640 0.9640 0.9640 0.9640 0.9640 0.9645 0.9640 0.9640 0.9640 0.9640
KNN 0.9493 0.9480 0.9480 0.9405 0.9475 0.8853 0.9485 0.9158 0.9240 0.9463 0.9405 0.9378 0.9378
LGBM 0.9640 0.9640 0.9640 0.9637 0.9637 0.9637 0.9637 0.9637 0.9640 0.9640 0.9640 0.9640 0.9637
LR 0.7627 0.7523 0.7523 0.7762 0.7770 0.7757 0.7775 0.7765 0.7497 0.5523 0.3238 0.7547 0.7680
MLP 0.9367 0.9280 0.9280 0.9502 0.9487 0.9448 0.9527 0.9545 0.9308 0.8677 0.5888 0.9242 0.9538
NB 0.6432 0.6432 0.6432 0.6432 0.6432 0.6432 0.6432 0.6432 0.6288 0.6432 0.6432 0.6303 0.6303
RF 0.9577 0.9577 0.9577 0.9570 0.9587 0.9568 0.9585 0.9580 0.9577 0.9582 0.9585 0.9587 0.9587
SVM 0.8135 0.8208 0.8208 0.8488 0.8333 0.7587 0.8090 0.8488 0.8193 0.4952 0.0522 0.8268 0.8415
TabNet 0.8963 0.8993 0.8993 0.8987 0.8997 0.8970 0.8942 0.8968 0.8827 0.9027 0.8840 0.8888 0.8907
XGBoost 0.9585 0.9585 0.9585 0.9585 0.9585 0.9585 0.9585 0.9585 0.9585 0.9585 0.9585 0.9585 0.9585
Magic Gamma Telescope Ada 0.8375 0.8375 0.8375 0.8375 0.8375 0.8375 0.8375 0.8375 0.8375 0.8375 0.8375 0.8375 0.8375
CART 0.8153 0.8153 0.8151 0.8153 0.8153 0.8134 0.8151 0.8151 0.8156 0.8134 0.8146 0.8151 0.8151
CatBoost 0.8891 0.8891 0.8891 0.8891 0.8891 0.8880 0.8891 0.8891 0.8891 0.8891 0.8889 0.8891 0.8891
KNN 0.8098 0.8332 0.8254 0.8340 0.8146 0.8393 0.8098 0.8340 0.8453 0.8091 0.8340 0.8419 0.8419
LGBM 0.8803 0.8785 0.8803 0.8792 0.8792 0.8808 0.8792 0.8810 0.8812 0.8803 0.8792 0.8792 0.8792
LR 0.7834 0.7914 0.7911 0.7936 0.7927 0.7939 0.7920 0.7936 0.8328 0.7893 0.6525 0.7976 0.7971
MLP 0.8170 0.8700 0.8677 0.8717 0.8722 0.8652 0.8517 0.8777 0.8729 0.8687 0.7730 0.8712 0.8777
NB 0.7275 0.7275 0.7275 0.7275 0.7275 0.7275 0.7275 0.7275 0.7934 0.7275 0.7273 0.7296 0.7296
RF 0.8808 0.8808 0.8808 0.8808 0.8808 0.8785 0.8808 0.8808 0.8803 0.8791 0.8819 0.8812 0.8812
SVM 0.2976 0.7483 0.5158 0.4341 0.4478 0.6851 0.3312 0.3212 0.3640 0.7536 0.3481 0.3467 0.3507
TabNet 0.8633 0.8757 0.8721 0.8701 0.8724 0.8731 0.8693 0.8791 0.8759 0.8789 0.8726 0.8731 0.8757
XGBoost 0.8803 0.8803 0.8803 0.8803 0.8803 0.8792 0.8803 0.8803 0.8831 0.8803 0.8835 0.8803 0.8803
Rice Cammeo And Osmancik Ada 0.9283 0.9283 0.9283 0.9283 0.9283 0.9283 0.9283 0.9283 0.9283 0.9283 0.9283 0.9283 0.9283
CART 0.8819 0.8819 0.8828 0.8828 0.8828 0.8828 0.8819 0.8819 0.8801 0.8819 0.8819 0.8819 0.8819
CatBoost 0.9265 0.9265 0.9265 0.9265 0.9265 0.9265 0.9265 0.9265 0.9265 0.9265 0.9274 0.9265 0.9265
KNN 0.8775 0.9143 0.9204 0.9143 0.9125 0.9108 0.8775 0.9064 0.9160 0.9151 0.9143 0.9134 0.9134
LGBM 0.9213 0.9221 0.9213 0.9178 0.9178 0.9178 0.9178 0.9160 0.9221 0.9213 0.9213 0.9213 0.9178
LR 0.9318 0.9335 0.9265 0.9300 0.9318 0.9318 0.9326 0.9300 0.9265 0.9151 0.5468 0.9300 0.9300
MLP 0.5468 0.9300 0.9335 0.9318 0.9239 0.9248 0.9116 0.9309 0.9265 0.9309 0.9169 0.9309 0.9326
NB 0.9151 0.9274 0.9274 0.9274 0.9274 0.9274 0.9151 0.9274 0.9204 0.9274 0.9274 0.9239 0.9239
RF 0.9265 0.9265 0.9265 0.9265 0.9265 0.9265 0.9265 0.9265 0.9248 0.9265 0.9265 0.9256 0.9256
SVM 0.9248 0.9291 0.9309 0.9309 0.8915 0.7970 0.8469 0.9274 0.9291 0.9274 0.8784 0.9274 0.9274
TabNet 0.5766 0.9248 0.7227 0.9335 0.9318 0.9265 0.9326 0.9265 0.9274 0.5494 0.4532 0.9283 0.9274
XGBoost 0.9125 0.9125 0.9125 0.9125 0.9125 0.9125 0.9125 0.9125 0.9125 0.9125 0.9125 0.9125 0.9125
Wine Ada 0.9259 0.9259 0.9259 0.9259 0.9259 0.9259 0.9259 0.9259 0.9259 0.9259 0.9259 0.9259 0.9259
CART 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630
CatBoost 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815
KNN 0.7407 0.9444 0.9444 0.9630 0.9444 0.9815 0.7407 0.9444 0.9630 0.8704 0.9630 0.9630 0.9630
LGBM 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815 0.9815
LR 0.9815 1.0000 1.0000 0.9815 1.0000 0.9815 1.0000 0.9815 1.0000 0.8704 0.3889 1.0000 1.0000
MLP 0.9815 0.9815 1.0000 0.9815 0.9815 0.9815 1.0000 0.9815 0.9815 1.0000 0.3889 0.9815 0.9815
NB 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 0.9815 0.9815
RF 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
SVM 0.5926 1.0000 1.0000 0.9815 0.9815 0.9815 0.6111 0.9815 0.9815 0.8148 0.3889 0.9815 0.9815
TabNet 0.2222 0.4074 0.2778 0.0926 0.1481 0.1852 0.2778 0.2037 0.2963 0.3333 0.2593 0.3704 0.2037
XGBoost 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630 0.9630
TABLE XII: Accuracy by dataset, model, and scaling method.
TABLE XIII: Time to Train (s) by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
Breast Cancer Wisconsin Diagnostic Ada 0.0625 0.0660 0.0623 0.0626 0.0661 0.0619 0.0618 0.0617 0.0616 0.0618 0.0621 0.0619 0.0618
CART 0.0035 0.0036 0.0039 0.0035 0.0040 0.0036 0.0035 0.0035 0.0035 0.0035 0.0036 0.0036 0.0036
CatBoost 0.9431 0.9772 0.9543 0.9512 0.9513 0.9340 0.9568 0.9348 0.9510 0.9467 0.9513 0.9505 0.9479
KNN 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0003
LGBM 0.0213 0.0306 0.0219 0.0275 0.0237 0.0233 0.0228 0.0256 0.0213 0.0210 0.0216 0.0213 0.0239
LR 0.0043 0.0047 0.0014 0.0012 0.0025 0.0024 0.0044 0.0012 0.0012 0.0011 0.0013 0.0012 0.0010
MLP 0.1708 0.2342 0.2929 0.1683 0.2549 0.1412 0.0107 0.1824 0.3734 0.3249 0.0070 0.3637 0.1990
NB 0.0003 0.0003 0.0003 0.0003 0.0004 0.0003 0.0004 0.0003 0.0003 0.0004 0.0003 0.0004 0.0004
RF 0.0706 0.0729 0.0707 0.0709 0.0711 0.0714 0.0709 0.0706 0.0704 0.0708 0.0706 0.0708 0.0708
SVM 0.0016 0.0008 0.0007 0.0009 0.0012 0.0014 0.0023 0.0009 0.0009 0.0010 0.0017 0.0008 0.0009
TabNet 0.0350 0.9545 0.0328 0.0314 0.0325 0.0339 0.0346 0.0319 0.0328 0.0344 0.0330 0.0349 0.0339
XGBoost 0.0151 0.0995 0.0155 0.0157 0.0153 0.0153 0.0151 0.0158 0.0154 0.0158 0.0152 0.0152 0.0159
Dry Bean Ada 0.7061 0.7334 0.7174 0.7231 0.7148 0.7192 0.7052 0.7145 0.7136 0.7101 0.7053 0.7134 0.7127
CART 0.1292 0.1307 0.1309 0.1312 0.1313 0.1304 0.1293 0.1302 0.1304 0.1298 0.1293 0.1304 0.1302
CatBoost 3.6706 3.6325 3.6400 3.6436 3.6461 3.6497 3.6579 3.6701 3.6637 3.6749 3.6736 3.6620 3.6708
KNN 0.0008 0.0008 0.0008 0.0008 0.0008 0.0008 0.0008 0.0008 0.0009 0.0010 0.0010 0.0009 0.0009
LGBM 0.2732 0.2741 0.2759 0.2988 0.2991 0.2989 0.2970 0.3025 0.2760 0.2743 0.2764 0.2748 0.2963
LR 0.1024 0.1288 0.0976 0.0783 0.0982 0.1048 0.1042 0.0820 0.0947 0.1013 0.0604 0.0972 0.0528
MLP 0.3396 2.8338 3.2012 2.4228 0.9144 0.5234 0.5023 3.8822 3.8263 2.6816 8.1463 2.8726 3.1601
NB 0.0012 0.0013 0.0012 0.0013 0.0013 0.0013 0.0013 0.0012 0.0012 0.0013 0.0013 0.0013 0.0013
RF 1.8703 1.9180 1.8860 1.8872 1.9258 1.8776 1.8653 1.8734 1.8702 2.0044 1.8541 1.8794 1.8811
SVM 0.1096 0.1588 0.1932 0.1560 0.1382 0.1516 0.0975 0.1470 0.1399 0.2449 1.0143 0.1336 0.1299
TabNet 11.8293 11.8382 11.7919 11.5408 11.7602 11.7540 11.3646 11.3732 11.5892 11.5038 11.7965 11.4746 11.4619
XGBoost 0.2849 0.2832 0.2870 0.2877 0.2877 0.2879 0.2882 0.2915 0.2901 0.2907 0.2895 0.2899 0.2963
Glass Identification Ada 0.0269 0.0269 0.0275 0.0291 0.0267 0.0267 0.0270 0.0265 0.0265 0.0268 0.0266 0.0266 0.0270
CART 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0006 0.0006 0.0007 0.0007 0.0007 0.0007
CatBoost 0.7934 0.7912 0.7935 0.8024 0.7970 0.8090 0.7974 0.8031 0.8038 0.7896 0.8072 0.8029 0.7971
KNN 0.0002 0.0003 0.0002 0.0002 0.0003 0.0003 0.0002 0.0004 0.0002 0.0003 0.0003 0.0003 0.0003
LGBM 0.0422 0.0433 0.0427 0.0345 0.0366 0.0348 0.0360 0.0410 0.0406 0.0400 0.0362 0.0362 0.0352
LR 0.0067 0.0023 0.0030 0.0020 0.0017 0.0072 0.0023 0.0026 0.0026 0.0021 0.0018 0.0029 0.0016
MLP 0.2307 0.2427 0.2416 0.2376 0.2341 0.2284 0.2333 0.2372 0.2360 0.2328 0.0195 0.2345 0.2357
NB 0.0003 0.0004 0.0005 0.0003 0.0004 0.0004 0.0004 0.0003 0.0003 0.0003 0.0004 0.0004 0.0005
RF 0.0453 0.0456 0.0456 0.0453 0.0482 0.0454 0.0453 0.0453 0.0453 0.0453 0.0455 0.0453 0.0462
SVM 0.0010 0.0007 0.0007 0.0008 0.0010 0.0022 0.0010 0.0009 0.0007 0.0007 0.0007 0.0007 0.0008
TabNet 0.0293 0.0338 0.0316 0.0301 0.0356 0.0317 0.0299 0.0293 0.0314 0.0297 0.0322 0.0310 0.0305
XGBoost 0.0332 0.0337 0.0341 0.0340 0.0339 0.0342 0.0334 0.0344 0.0340 0.0340 0.0339 0.0357 0.0354
Heart Disease Ada 0.0265 0.0266 0.0272 0.0272 0.0265 0.0266 0.0265 0.0264 0.0264 0.0266 0.0264 0.0265 0.0269
CART 0.0007 0.0007 0.0008 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007
CatBoost 0.6742 0.6528 0.6595 0.6623 0.6676 0.6681 0.6694 0.6754 0.6734 0.6811 0.6725 0.6683 0.6687
KNN 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
LGBM 0.0466 0.0433 0.0450 0.0412 0.0405 0.0402 0.0390 0.0439 0.0418 0.0449 0.0385 0.0388 0.0391
LR 0.0073 0.0044 0.0054 0.0018 0.0058 0.0069 0.0078 0.0031 0.0046 0.0036 0.0024 0.0045 0.0016
MLP 0.0198 0.1104 0.1105 0.3423 0.2924 0.3025 0.0822 0.3942 0.3946 0.1659 0.0234 0.2356 0.3996
NB 0.0003 0.0003 0.0004 0.0003 0.0003 0.0003 0.0003 0.0003 0.0004 0.0003 0.0004 0.0005 0.0005
RF 0.0449 0.0453 0.0452 0.0451 0.0481 0.0453 0.0453 0.0451 0.0451 0.0451 0.0451 0.0450 0.0476
SVM 0.0029 0.0009 0.0009 0.0017 0.0028 0.0031 0.0030 0.0012 0.0010 0.0010 0.0008 0.0009 0.0011
TabNet 0.0291 0.0330 0.0329 0.1327 0.0333 0.0334 0.0305 0.0305 0.0307 0.0301 0.0324 0.0306 0.0307
XGBoost 0.0379 0.0388 0.0387 0.0391 0.0387 0.0391 0.0382 0.0396 0.0388 0.0388 0.0386 0.0404 0.0407
Iris Ada 0.0238 0.0239 0.0246 0.0243 0.0236 0.0258 0.0237 0.0234 0.0235 0.0236 0.0236 0.0236 0.0239
CART 0.0003 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0003 0.0003 0.0003 0.0004 0.0004 0.0003
CatBoost 0.2805 0.2828 0.2800 0.2777 0.2863 0.2898 0.2740 0.2748 0.2893 0.2820 0.2874 0.2870 0.2846
KNN 0.0004 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0003 0.0003 0.0002 0.0002 0.0002 0.0003
LGBM 0.0116 0.0126 0.0119 0.0130 0.0136 0.0127 0.0126 0.0118 0.0121 0.0116 0.0123 0.0122 0.0115
LR 0.0045 0.0014 0.0018 0.0014 0.0013 0.0044 0.0015 0.0015 0.0014 0.0015 0.0012 0.0013 0.0012
MLP 0.1074 0.1412 0.1444 0.0947 0.0969 0.1067 0.0921 0.1164 0.1309 0.1910 0.0049 0.1556 0.1173
NB 0.0003 0.0003 0.0004 0.0003 0.0004 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
RF 0.0382 0.0387 0.0387 0.0386 0.0399 0.0385 0.0385 0.0385 0.0384 0.0383 0.0385 0.0384 0.0399
SVM 0.0003 0.0003 0.0004 0.0003 0.0004 0.0004 0.0003 0.0003 0.0004 0.0004 0.0004 0.0004 0.0004
TabNet 0.0285 0.0310 0.0322 0.0282 0.0320 0.0306 0.0304 0.0296 0.0299 0.0290 0.0301 0.0307 0.0296
XGBoost 0.0130 0.0141 0.0136 0.0134 0.0134 0.0135 0.0132 0.0136 0.0134 0.0135 0.0135 0.0142 0.0144
Letter Recognition Ada 0.3581 0.3601 0.3589 0.3570 0.3563 0.3565 0.3555 0.3551 0.3548 0.3554 0.3550 0.3603 0.3546
CART 0.0381 0.0380 0.0382 0.0381 0.0378 0.0384 0.0381 0.0378 0.0379 0.0380 0.0380 0.0379 0.0380
CatBoost 5.8619 5.8869 5.8750 5.8861 5.8526 5.8972 5.8645 5.8823 5.8914 5.8835 5.8919 5.8794 5.8696
KNN 0.0010 0.0010 0.0009 0.0009 0.0009 0.0009 0.0009 0.0010 0.0011 0.0012 0.0011 0.0011 0.0011
LGBM 1.1319 1.1009 1.0970 1.1101 1.1199 1.1223 1.1144 1.1380 1.0928 1.1109 1.1247 1.0959 1.1216
LR 0.2923 0.3110 0.3040 0.1714 0.2318 0.2898 0.2954 0.2229 0.2969 0.0625 0.0764 0.2906 0.1379
MLP 11.1298 22.3418 22.0963 11.5231 10.0211 8.8762 8.8380 13.5357 22.1324 22.2006 18.4904 22.0306 14.9389
NB 0.0029 0.0026 0.0026 0.0025 0.0026 0.0025 0.0026 0.0025 0.0024 0.0026 0.0025 0.0025 0.0025
RF 0.8526 0.8806 0.8582 0.8632 0.8850 0.8594 0.9108 0.8535 0.8573 0.8554 0.8554 0.8547 0.8557
SVM 0.8484 0.7881 0.7841 0.7588 0.8009 0.9311 0.8348 0.7936 0.6644 2.6349 3.5273 0.6776 0.6531
TabNet 16.6080 16.2526 16.6332 16.3683 16.3003 16.4503 16.9526 16.6139 16.5979 16.4259 16.5634 16.2669 16.6276
XGBoost 0.7149 0.6430 0.6358 0.6357 0.6335 0.6344 0.6589 0.6445 0.6438 0.6458 0.7151 0.6552 0.6772
Magic Gamma Telescope Ada 0.5774 0.5811 0.5844 0.5766 0.5774 0.5807 0.5775 0.5764 0.5760 0.5751 0.5624 0.5781 0.5763
CART 0.1448 0.1449 0.1456 0.1454 0.1448 0.1450 0.1451 0.1448 0.1448 0.1443 0.1436 0.1450 0.1462
CatBoost 2.4467 2.4322 2.4213 2.4526 2.4666 2.4253 2.4503 2.4659 2.4766 2.4590 2.4641 2.4541 2.4565
KNN 0.0065 0.0065 0.0065 0.0065 0.0065 0.0065 0.0065 0.0071 0.0069 0.0075 0.0068 0.0069 0.0068
LGBM 0.0469 0.0457 0.0460 0.0461 0.0457 0.0461 0.0464 0.0468 0.0444 0.0468 0.0447 0.0447 0.0464
LR 0.0283 0.0074 0.0106 0.0063 0.0176 0.0124 0.0269 0.0067 0.0081 0.0089 0.0066 0.0068 0.0054
MLP 0.5592 4.3368 4.8919 4.0988 4.1299 4.0089 1.6349 3.4037 4.5621 7.1413 5.2403 3.0823 5.8520
NB 0.0013 0.0013 0.0015 0.0013 0.0013 0.0014 0.0013 0.0014 0.0013 0.0013 0.0016 0.0013 0.0014
RF 2.6338 2.6490 2.6361 2.6471 2.6767 2.6233 2.6346 2.6319 2.6293 2.6260 2.6381 2.6242 2.6470
SVM 0.0697 0.2825 0.2742 0.2290 0.0940 0.1469 0.0727 0.2240 0.2731 0.2712 0.2685 0.2738 0.2736
TabNet 16.3108 16.2803 16.3504 16.1869 17.0110 16.9746 16.7629 16.9795 16.5454 16.8754 16.2501 16.7170 16.5016
XGBoost 0.0501 0.0494 0.0496 0.0501 0.0495 0.0495 0.0495 0.0500 0.0491 0.0543 0.0499 0.0510 0.0496
Rice Cammeo And Osmancik Ada 0.0993 0.1006 0.1024 0.0991 0.0992 0.1018 0.0991 0.0992 0.0992 0.1008 0.0990 0.0995 0.0994
CART 0.0094 0.0104 0.0096 0.0103 0.0095 0.0094 0.0094 0.0094 0.0094 0.0095 0.0094 0.0094 0.0095
CatBoost 0.9131 0.8989 0.8965 0.9065 0.9136 0.9021 0.9150 0.9123 0.9078 0.9154 0.9173 0.9184 0.9123
KNN 0.0009 0.0010 0.0010 0.0010 0.0009 0.0009 0.0009 0.0010 0.0011 0.0012 0.0010 0.0010 0.0012
LGBM 0.0294 0.0295 0.0291 0.0309 0.0312 0.0307 0.0310 0.0315 0.0292 0.0296 0.0296 0.0298 0.0309
LR 0.0081 0.0016 0.0034 0.0018 0.0028 0.0055 0.0075 0.0017 0.0015 0.0032 0.0019 0.0018 0.0015
MLP 0.0464 0.2474 0.6241 0.1954 0.3662 0.1570 0.1310 0.2864 0.4593 1.5107 2.6258 0.2457 0.2491
NB 0.0004 0.0005 0.0005 0.0005 0.0005 0.0004 0.0004 0.0005 0.0005 0.0004 0.0004 0.0004 0.0005
RF 0.2023 0.2036 0.2032 0.2026 0.2086 0.2033 0.2024 0.2034 0.2023 0.2026 0.2022 0.2020 0.2106
SVM 0.0093 0.0173 0.0223 0.0201 0.0146 0.0104 0.0074 0.0199 0.0176 0.0361 0.0551 0.0168 0.0170
TabNet 2.5276 2.6327 2.6146 2.5252 2.5909 2.7084 2.5272 2.7104 2.5292 2.5641 2.6258 2.6799 2.5631
XGBoost 0.0262 0.0262 0.0262 0.0263 0.0263 0.0262 0.0260 0.0267 0.0263 0.0264 0.0259 0.0275 0.0260
Wine Ada 0.0271 0.0280 0.0276 0.0271 0.0272 0.0283 0.0277 0.0269 0.0270 0.0271 0.0272 0.0274 0.0274
CART 0.0005 0.0005 0.0005 0.0006 0.0006 0.0005 0.0006 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005
CatBoost 0.5668 0.5651 0.5670 0.5647 0.5677 0.5665 0.5729 0.5716 0.5675 0.5645 0.5724 0.5610 0.5577
KNN 0.0003 0.0002 0.0003 0.0002 0.0002 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
LGBM 0.0164 0.0149 0.0151 0.0167 0.0162 0.0144 0.0160 0.0154 0.0152 0.0144 0.0156 0.0175 0.0154
LR 0.0059 0.0015 0.0031 0.0014 0.0040 0.0031 0.0067 0.0013 0.0019 0.0013 0.0019 0.0019 0.0012
MLP 0.2105 0.1228 0.1658 0.0603 0.0784 0.0377 0.1086 0.0696 0.1100 0.2217 0.0076 0.1234 0.0789
NB 0.0003 0.0003 0.0003 0.0004 0.0004 0.0003 0.0004 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
RF 0.0406 0.0410 0.0408 0.0411 0.0411 0.0408 0.0410 0.0406 0.0406 0.0407 0.0407 0.0405 0.0410
SVM 0.0008 0.0004 0.0004 0.0005 0.0006 0.0005 0.0008 0.0005 0.0004 0.0005 0.0006 0.0005 0.0005
TabNet 0.0291 0.0298 0.0289 0.0286 0.0319 0.0315 0.0321 0.0307 0.0306 0.0304 0.0310 0.0336 0.0292
XGBoost 0.0122 0.0120 0.0121 0.0121 0.0119 0.0119 0.0119 0.0124 0.0121 0.0121 0.0118 0.0126 0.0118
TABLE XIII: Time to Train (s) by dataset, model, and scaling method.
TABLE XIV: Time to Inference (s) by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
Breast Cancer Wisconsin Diagnostic Ada 0.0018 0.0019 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018
CART 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
CatBoost 0.0005 0.0005 0.0006 0.0004 0.0005 0.0006 0.0005 0.0004 0.0004 0.0004 0.0005 0.0007 0.0006
KNN 0.0025 0.0131 0.0025 0.0025 0.0025 0.0025 0.0025 0.0030 0.0025 0.0025 0.0034 0.0027 0.0028
LGBM 0.0004 0.0006 0.0004 0.0007 0.0004 0.0004 0.0004 0.0007 0.0004 0.0004 0.0004 0.0005 0.0005
LR 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
MLP 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
NB 0.0001 0.0001 0.0001 0.0002 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
RF 0.0013 0.0014 0.0013 0.0013 0.0013 0.0013 0.0013 0.0013 0.0013 0.0013 0.0013 0.0013 0.0013
SVM 0.0002 0.0002 0.0002 0.0001 0.0001 0.0001 0.0003 0.0001 0.0002 0.0003 0.0006 0.0002 0.0002
TabNet 0.0037 0.0145 0.0042 0.0038 0.0037 0.0036 0.0040 0.0038 0.0040 0.0038 0.0037 0.0039 0.0042
XGBoost 0.0003 0.0005 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
Dry Bean Ada 0.0073 0.0074 0.0075 0.0078 0.0073 0.0073 0.0074 0.0072 0.0071 0.0072 0.0072 0.0073 0.0073
CART 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
CatBoost 0.0024 0.0027 0.0025 0.0029 0.0026 0.0022 0.0023 0.0022 0.0024 0.0025 0.0023 0.0022 0.0022
KNN 0.0529 0.0532 0.0536 0.0536 0.0527 0.0539 0.0530 0.0594 0.0553 0.0554 0.0542 0.0553 0.0565
LGBM 0.0101 0.0097 0.0099 0.0099 0.0100 0.0101 0.0099 0.0100 0.0099 0.0099 0.0100 0.0102 0.0098
LR 0.0002 0.0003 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002
MLP 0.0011 0.0012 0.0012 0.0012 0.0011 0.0011 0.0012 0.0012 0.0012 0.0010 0.0011 0.0010 0.0010
NB 0.0007 0.0009 0.0006 0.0007 0.0007 0.0008 0.0007 0.0006 0.0007 0.0007 0.0007 0.0007 0.0007
RF 0.0180 0.0182 0.0181 0.0184 0.0188 0.0180 0.0182 0.0179 0.0183 0.0182 0.0187 0.0187 0.0179
SVM 0.0255 0.1381 0.1723 0.0892 0.0597 0.0384 0.0138 0.0907 0.1141 0.2042 0.4215 0.1154 0.0968
TabNet 0.0237 0.0238 0.0238 0.0238 0.0243 0.0244 0.0239 0.0239 0.0238 0.0236 0.0243 0.0238 0.0236
XGBoost 0.0035 0.0035 0.0035 0.0035 0.0035 0.0035 0.0035 0.0035 0.0036 0.0036 0.0035 0.0036 0.0036
Glass Identification Ada 0.0017 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016
CART 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
CatBoost 0.0004 0.0006 0.0006 0.0004 0.0006 0.0005 0.0006 0.0005 0.0006 0.0008 0.0007 0.0005 0.0007
KNN 0.0012 0.0013 0.0012 0.0012 0.0012 0.0012 0.0012 0.0014 0.0014 0.0013 0.0014 0.0014 0.0014
LGBM 0.0005 0.0006 0.0006 0.0006 0.0006 0.0005 0.0005 0.0006 0.0005 0.0005 0.0006 0.0005 0.0005
LR 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
MLP 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
NB 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
RF 0.0013 0.0013 0.0013 0.0013 0.0013 0.0013 0.0013 0.0013 0.0012 0.0013 0.0013 0.0013 0.0013
SVM 0.0002 0.0002 0.0002 0.0002 0.0002 0.0001 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002
TabNet 0.0033 0.0037 0.0031 0.0036 0.0033 0.0034 0.0036 0.0033 0.0033 0.0034 0.0035 0.0034 0.0035
XGBoost 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0005 0.0004 0.0005 0.0004
Heart Disease Ada 0.0017 0.0016 0.0016 0.0016 0.0016 0.0017 0.0017 0.0016 0.0016 0.0017 0.0017 0.0016 0.0017
CART 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
CatBoost 0.0005 0.0005 0.0005 0.0006 0.0006 0.0006 0.0004 0.0008 0.0006 0.0006 0.0007 0.0005 0.0007
KNN 0.0016 0.0016 0.0016 0.0016 0.0015 0.0016 0.0016 0.0020 0.0018 0.0017 0.0016 0.0019 0.0017
LGBM 0.0007 0.0006 0.0006 0.0006 0.0006 0.0007 0.0005 0.0006 0.0006 0.0005 0.0006 0.0006 0.0005
LR 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
MLP 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
NB 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0002 0.0001 0.0001 0.0001
RF 0.0014 0.0014 0.0014 0.0014 0.0015 0.0014 0.0014 0.0015 0.0014 0.0014 0.0015 0.0014 0.0014
SVM 0.0003 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002
TabNet 0.0034 0.0036 0.0035 0.0036 0.0038 0.0035 0.0036 0.0036 0.0035 0.0035 0.0033 0.0034 0.0035
XGBoost 0.0005 0.0006 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0006 0.0005
Iris Ada 0.0015 0.0015 0.0016 0.0015 0.0016 0.0015 0.0015 0.0016 0.0016 0.0015 0.0016 0.0015 0.0016
CART 0.0000 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0000 0.0000 0.0001 0.0001 0.0001
CatBoost 0.0006 0.0005 0.0005 0.0005 0.0005 0.0006 0.0004 0.0004 0.0004 0.0005 0.0005 0.0005 0.0004
KNN 0.0012 0.0010 0.0010 0.0009 0.0009 0.0009 0.0009 0.0012 0.0010 0.0010 0.0010 0.0011 0.0010
LGBM 0.0004 0.0004 0.0005 0.0004 0.0004 0.0006 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004
LR 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0000 0.0001 0.0001 0.0001 0.0001
MLP 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
NB 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
RF 0.0011 0.0011 0.0011 0.0011 0.0011 0.0011 0.0011 0.0011 0.0011 0.0011 0.0011 0.0011 0.0011
SVM 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
TabNet 0.0033 0.0036 0.0034 0.0034 0.0032 0.0034 0.0032 0.0034 0.0033 0.0032 0.0035 0.0032 0.0035
XGBoost 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
Letter Recognition Ada 0.0204 0.0200 0.0200 0.0193 0.0200 0.0197 0.0195 0.0194 0.0197 0.0198 0.0197 0.0197 0.0196
CART 0.0006 0.0006 0.0006 0.0006 0.0006 0.0006 0.0006 0.0006 0.0006 0.0006 0.0006 0.0006 0.0006
CatBoost 0.0049 0.0049 0.0053 0.0044 0.0052 0.0044 0.0046 0.0050 0.0050 0.0045 0.0050 0.0048 0.0044
KNN 0.8666 0.0829 0.0819 0.0821 0.0823 0.0828 0.0825 0.0870 0.0841 0.0903 0.0836 0.0854 0.0854
LGBM 0.0611 0.0606 0.0606 0.0610 0.0613 0.0610 0.0612 0.0601 0.0607 0.0609 0.0606 0.0609 0.0609
LR 0.0006 0.0006 0.0005 0.0005 0.0005 0.0005 0.0006 0.0005 0.0005 0.0005 0.0004 0.0005 0.0005
MLP 0.0022 0.0022 0.0022 0.0021 0.0022 0.0021 0.0021 0.0021 0.0022 0.0022 0.0022 0.0022 0.0022
NB 0.0036 0.0032 0.0031 0.0033 0.0031 0.0031 0.0030 0.0030 0.0029 0.0030 0.0030 0.0030 0.0029
RF 0.0471 0.0480 0.0473 0.0472 0.0493 0.0481 0.0484 0.0479 0.0483 0.0482 0.0480 0.0482 0.0486
SVM 0.8372 1.3641 1.3598 0.8540 0.8260 0.8683 0.8521 0.8722 1.1584 1.8609 1.8765 1.2526 1.0256
TabNet 0.0354 0.0355 0.0357 0.0352 0.0351 0.0354 0.0354 0.0366 0.0356 0.0358 0.0357 0.0354 0.0351
XGBoost 0.0206 0.0208 0.0207 0.0207 0.0208 0.0208 0.0209 0.0208 0.0210 0.0205 0.0206 0.0207 0.0211
Magic Gamma Telescope Ada 0.0073 0.0071 0.0071 0.0068 0.0068 0.0071 0.0069 0.0067 0.0067 0.0068 0.0067 0.0068 0.0068
CART 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0004 0.0005 0.0004 0.0005 0.0005
CatBoost 0.0016 0.0019 0.0019 0.0017 0.0021 0.0019 0.0019 0.0018 0.0017 0.0017 0.0020 0.0019 0.0019
KNN 0.1088 0.1324 0.1550 0.1659 0.1254 0.1007 0.1093 0.1635 0.1617 0.1156 0.1673 0.1543 0.1536
LGBM 0.0024 0.0025 0.0023 0.0024 0.0022 0.0022 0.0023 0.0023 0.0024 0.0023 0.0023 0.0023 0.0023
LR 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
MLP 0.0012 0.0012 0.0012 0.0012 0.0012 0.0012 0.0012 0.0012 0.0011 0.0011 0.0011 0.0010 0.0011
NB 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0002 0.0003 0.0003 0.0002 0.0003 0.0003 0.0003
RF 0.0347 0.0344 0.0349 0.0349 0.0341 0.0346 0.0342 0.0341 0.0341 0.0341 0.0343 0.0344 0.0345
SVM 0.0214 0.1030 0.1027 0.0838 0.0291 0.0523 0.0232 0.0856 0.1028 0.1025 0.1023 0.1031 0.1024
TabNet 0.0342 0.0336 0.0340 0.0339 0.0352 0.0337 0.0342 0.0351 0.0339 0.0340 0.0341 0.0349 0.0339
XGBoost 0.0009 0.0009 0.0011 0.0010 0.0009 0.0009 0.0009 0.0009 0.0009 0.0010 0.0009 0.0009 0.0009
Rice Cammeo And Osmancik Ada 0.0026 0.0026 0.0026 0.0026 0.0026 0.0027 0.0025 0.0026 0.0026 0.0026 0.0026 0.0025 0.0026
CART 0.0001 0.0001 0.0001 0.0002 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
CatBoost 0.0009 0.0011 0.0009 0.0009 0.0009 0.0008 0.0007 0.0009 0.0010 0.0009 0.0007 0.0007 0.0010
KNN 0.0139 0.0151 0.0148 0.0149 0.0141 0.0147 0.0140 0.0152 0.0153 0.0172 0.0155 0.0155 0.0154
LGBM 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0008 0.0007 0.0007 0.0007 0.0007 0.0007
LR 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
MLP 0.0003 0.0003 0.0003 0.0002 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0002
NB 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
RF 0.0045 0.0043 0.0044 0.0045 0.0045 0.0044 0.0044 0.0044 0.0043 0.0044 0.0044 0.0044 0.0044
SVM 0.0012 0.0061 0.0083 0.0055 0.0031 0.0017 0.0007 0.0056 0.0058 0.0140 0.0213 0.0058 0.0055
TabNet 0.0090 0.0091 0.0090 0.0091 0.0090 0.0090 0.0089 0.0091 0.0090 0.0090 0.0095 0.0099 0.0089
XGBoost 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005
Wine Ada 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016 0.0016 0.0015 0.0015
CART 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
CatBoost 0.0005 0.0005 0.0005 0.0004 0.0004 0.0007 0.0006 0.0007 0.0005 0.0005 0.0005 0.0005 0.0005
KNN 0.0011 0.0011 0.0011 0.0011 0.0011 0.0011 0.0011 0.0011 0.0012 0.0011 0.0012 0.0012 0.0012
LGBM 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0005 0.0007 0.0004
LR 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
MLP 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
NB 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
RF 0.0011 0.0011 0.0011 0.0011 0.0011 0.0011 0.0012 0.0011 0.0011 0.0011 0.0011 0.0012 0.0011
SVM 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
TabNet 0.0034 0.0034 0.0033 0.0033 0.0031 0.0032 0.0036 0.0035 0.0033 0.0036 0.0037 0.0035 0.0032
XGBoost 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
TABLE XIV: Time to Inference (s) by dataset, model, and scaling method.
TABLE XV: $R^{2}$ Score by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
Abalone Ada 0.1671 0.2210 0.2243 0.2503 0.1510 0.1938 0.1296 0.2301 0.1347 0.2018 0.1690 0.2112 0.2112
CART 0.1383 0.1464 0.1477 0.1560 0.1456 0.1500 0.1626 0.1417 0.1663 0.1416 0.1373 0.1686 0.1686
CatBoost 0.5226 0.5226 0.5225 0.5226 0.5225 0.5225 0.5225 0.5227 0.5224 0.5226 0.5226 0.5223 0.5223
KNN 0.5164 0.5023 0.4955 0.4662 0.5056 0.4406 0.5164 0.4552 0.4766 0.4100 0.4662 0.4699 0.4699
LGBM 0.5260 0.5278 0.5260 0.5256 0.5256 0.5256 0.5256 0.5190 0.5277 0.5260 0.5257 0.5257 0.5259
LinearRegression 0.5150 0.5150 0.5150 0.5150 0.5150 0.5150 0.5150 0.5150 0.4964 0.5150 0.5151 0.5204 0.5204
MLP 0.5245 0.5276 0.5265 0.5578 0.5580 0.5443 0.5559 0.5632 0.5434 0.5300 0.0226 0.5337 0.5560
RF 0.5244 0.5247 0.5249 0.5234 0.5244 0.5248 0.5247 0.5241 0.5230 0.5232 0.5239 0.5236 0.5236
SVR 0.5293 0.5283 0.5257 0.5421 0.5505 0.5157 0.5430 0.5398 0.5345 0.3615 0.5421 0.5437 0.5437
TabNet 0.3127 0.3609 0.2115 0.5131 0.5224 0.5554 0.3745 0.5434 0.4144 -0.0595 -0.0462 0.3457 0.5017
XGBoost 0.4546 0.4546 0.4546 0.4546 0.4546 0.4546 0.4546 0.4546 0.4546 0.4546 0.4546 0.4546 0.4546
Air Quality Ada 0.9992 0.9992 0.9984 0.9984 0.9984 0.9984 0.9992 0.9992 0.9984 0.9989 0.9984 0.9984 0.9984
CART 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
CatBoost 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999
KNN 0.9995 0.9993 0.9993 0.9994 0.9994 0.9996 0.9995 0.9986 0.9986 0.9993 0.9994 0.9994 0.9994
LGBM 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999
LinearRegression 0.9992 0.9992 0.9992 0.9992 0.9992 0.9992 0.9992 0.9992 0.8736 0.9992 0.9992 0.9977 0.9977
MLP 0.9985 0.9990 1.0000 1.0000 0.9999 1.0000 0.9975 0.9999 0.9983 0.9992 0.8187 0.9997 0.9999
RF 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
SVR 0.9966 0.9792 0.9619 0.9269 0.9435 0.9915 0.9675 0.9188 0.7749 0.9540 0.9269 0.9297 0.9297
TabNet 0.9997 0.9984 0.9995 0.9978 0.9998 0.9981 0.9993 0.9998 0.9948 0.9869 0.0353 0.9998 0.9990
XGBoost 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999 0.9999
Appliances Energy Prediction Ada -2.6328 -0.7955 -3.8255 -3.4419 -3.9861 -3.6762 -1.4564 -1.4564 -1.7618 -0.7955 -1.4562 -0.7955 -0.7955
CART 0.1358 0.1348 0.1361 0.1378 0.1378 0.1367 0.1354 0.1377 0.1540 0.1406 0.1398 0.1410 0.1410
CatBoost 0.4529 0.4570 0.4529 0.4570 0.4570 0.4570 0.4570 0.4570 0.4588 0.4529 0.4568 0.4570 0.4570
KNN 0.1681 0.2875 0.2049 0.3279 0.2595 0.4720 0.1681 0.2929 0.3228 0.1568 0.3279 0.3280 0.3280
LGBM 0.4318 0.4276 0.4318 0.4334 0.4334 0.4334 0.4334 0.4192 0.4285 0.4318 0.4351 0.4351 0.4334
LinearRegression 0.1672 0.1672 0.1672 0.1672 0.1672 0.1672 0.1672 0.1672 0.1488 0.1672 0.1672 0.1644 0.1644
MLP 0.1598 0.2037 0.1581 0.3144 0.3093 0.3161 0.3017 0.2970 0.2243 0.1540 0.0003 0.2209 0.2936
RF 0.5122 0.5122 0.5120 0.5122 0.5129 0.5124 0.5127 0.5125 0.5114 0.5132 0.5120 0.5126 0.5126
SVR -0.1056 -0.0032 -0.0275 0.0154 0.0041 -0.0518 -0.0152 -0.0096 0.0101 -0.0409 0.0154 0.0110 0.0110
TabNet 0.3054 0.2955 0.3177 0.2828 0.3031 0.3333 0.3049 0.2775 0.2997 0.2838 0.3258 0.2973 0.3195
XGBoost 0.4788 0.4788 0.4788 0.4788 0.4788 0.4788 0.4788 0.4788 0.4841 0.4788 0.4788 0.4788 0.4788
Concrete Compressive Strength Ada 0.7762 0.7761 0.7717 0.7782 0.7783 0.7788 0.7829 0.7788 0.7797 0.7788 0.7814 0.7874 0.7874
CART 0.8252 0.8252 0.8252 0.8252 0.8252 0.8252 0.8251 0.8251 0.8186 0.8252 0.8251 0.8250 0.8250
CatBoost 0.9337 0.9337 0.9337 0.9337 0.9337 0.9338 0.9337 0.9338 0.9344 0.9338 0.9337 0.9338 0.9338
KNN 0.6770 0.6472 0.6631 0.6714 0.7057 0.4853 0.6773 0.7446 0.7971 0.6381 0.6714 0.7119 0.7119
LGBM 0.9229 0.9229 0.9229 0.9217 0.9217 0.9217 0.9217 0.9226 0.9234 0.9229 0.9232 0.9232 0.9219
LinearRegression 0.5944 0.5944 0.5944 0.5944 0.5944 0.5944 0.5944 0.5944 0.8056 0.5944 0.5945 0.6983 0.6983
MLP 0.8030 0.7922 0.7468 0.8725 0.8744 0.8319 0.8007 0.8662 0.8333 0.5683 -0.0024 0.7430 0.8341
RF 0.8896 0.8895 0.8895 0.8891 0.8894 0.8895 0.8893 0.8894 0.8904 0.8895 0.8895 0.8897 0.8897
SVR 0.2259 0.5593 0.5394 0.6093 0.5897 0.3754 0.5446 0.6987 0.7302 0.3923 0.6093 0.6315 0.6315
TabNet -8297.0175 -4.5391 -4.5299 -4.3532 -7.9591 -3.8492 -876.1911 -4.4619 -4.5697 -4.5216 -4.6802 -4.6035 -4.4647
XGBoost 0.9104 0.9104 0.9104 0.9104 0.9104 0.9104 0.9104 0.9104 0.9104 0.9104 0.9104 0.9104 0.9104
Forest Fires Ada -0.4980 -0.4851 -0.4430 -0.0558 -0.0628 -0.4720 -0.4738 -0.4412 -0.0462 -0.4570 -0.4833 -0.0191 -0.0191
CART -0.6686 -0.6694 -0.6694 -0.6694 -0.6653 -0.6686 -0.6694 -0.6694 -0.6432 -0.6429 -0.6694 -0.6423 -0.6423
CatBoost -0.0437 -0.0437 -0.0437 -0.0437 -0.0437 -0.0437 -0.0437 -0.0437 -0.0438 -0.0437 -0.0437 -0.0437 -0.0437
KNN -0.0115 -0.0619 -0.0470 -0.0447 0.0345 -0.0530 -0.0115 -0.0345 0.0105 -0.0548 -0.0447 -0.0403 -0.0403
LGBM -0.0246 -0.0246 -0.0246 -0.0188 -0.0188 -0.0188 -0.0188 -0.0135 -0.0288 -0.0246 -0.0181 -0.0181 -0.0127
LinearRegression 0.0041 0.0041 0.0041 0.0041 0.0041 0.0041 0.0041 0.0041 0.0092 0.0041 0.0041 0.0061 0.0061
MLP -0.0067 0.0069 0.0057 0.0013 -0.0266 -0.0007 -0.0479 0.0121 0.0134 0.0062 -0.0012 0.0089 0.0144
RF -0.1060 -0.1084 -0.1101 -0.1093 -0.1093 -0.0995 -0.1094 -0.1058 -0.1004 -0.1102 -0.1060 -0.1060 -0.1060
SVR -0.0257 -0.0246 -0.0246 -0.0244 -0.0256 -0.0248 -0.0265 -0.0245 -0.0249 -0.0252 -0.0244 -0.0242 -0.0242
TabNet -1.1851 -0.0274 -0.0264 -0.0256 -0.0435 -0.0301 -1.2978 -0.0263 -0.0268 -0.0275 -0.0264 -0.0260 -0.0272
XGBoost -0.4684 -0.4684 -0.4684 -0.4684 -0.4684 -0.4684 -0.4684 -0.4684 -0.4684 -0.4684 -0.4684 -0.4684 -0.4684
Real Estate Valuation Ada 0.6579 0.6843 0.6843 0.6816 0.6843 0.6719 0.6726 0.6643 0.6842 0.6980 0.6678 0.6829 0.6829
CART 0.5608 0.5623 0.5623 0.5623 0.5607 0.5639 0.5607 0.5608 0.5567 0.6011 0.5623 0.5567 0.5567
CatBoost 0.7324 0.7325 0.7325 0.7324 0.7325 0.7325 0.7327 0.7325 0.7301 0.7325 0.7327 0.7322 0.7322
KNN 0.6232 0.6191 0.6232 0.6153 0.6488 0.4689 0.6232 0.6348 0.6924 0.5645 0.6153 0.6253 0.6253
LGBM 0.7001 0.7001 0.7001 0.7120 0.7120 0.7120 0.7120 0.7075 0.7002 0.7001 0.7000 0.7000 0.7120
LinearRegression 0.5601 0.5601 0.5601 0.5601 0.5601 0.5601 0.5601 0.5601 0.6805 0.5601 0.5601 0.6052 0.6052
MLP 0.6199 0.5762 0.5608 0.6436 0.6641 -4.7466 0.2915 0.6894 0.6388 0.5184 -0.0077 0.5586 0.6535
RF 0.7444 0.7449 0.7449 0.7444 0.7438 0.7450 0.7444 0.7444 0.7518 0.7456 0.7449 0.7430 0.7430
SVR 0.4897 0.5610 0.5327 0.5788 0.5481 0.4095 0.5202 0.5872 0.6311 0.5200 0.5788 0.5948 0.5948
TabNet -23524.2610 -8.6955 -8.8303 -8.6971 -8.4680 -2124512.8319 -1294.4512 -8.4379 -8.6292 -8.2463 -8.5425 -8.6645 -8.2867
XGBoost 0.7013 0.7013 0.7013 0.7013 0.7013 0.7013 0.7013 0.7013 0.7013 0.7013 0.7013 0.7013 0.7013
Wine Quality Ada 0.3092 0.3142 0.3013 0.3017 0.2968 0.2998 0.2994 0.3158 0.2956 0.3021 0.3148 0.3148 0.3148
CART 0.0308 0.0231 0.0273 0.0287 0.0238 0.0371 0.0357 0.0322 0.0118 0.0238 0.0280 0.0202 0.0202
CatBoost 0.4707 0.4708 0.4707 0.4707 0.4707 0.4708 0.4707 0.4707 0.4708 0.4706 0.4707 0.4710 0.4710
KNN 0.1221 0.2954 0.3283 0.3475 0.2706 0.1479 0.1221 0.3234 0.3245 0.3132 0.3473 0.3396 0.3396
LGBM 0.4557 0.4557 0.4557 0.4578 0.4578 0.4578 0.4578 0.4602 0.4556 0.4557 0.4555 0.4555 0.4578
LinearRegression 0.2701 0.2701 0.2701 0.2701 0.2701 0.2701 0.2701 0.2701 0.2723 0.2701 0.2702 0.2811 0.2811
MLP 0.2500 0.3455 0.3243 0.3872 0.3713 0.2518 0.3026 0.3879 0.3553 0.2928 -0.0026 0.3344 0.3864
RF 0.4985 0.4987 0.4991 0.4982 0.4994 0.4986 0.4990 0.4992 0.4986 0.4980 0.4979 0.4993 0.4993
SVR 0.1573 0.3561 0.3185 0.3842 0.3320 0.2115 0.2286 0.3799 0.3825 0.3128 0.3842 0.3907 0.3907
TabNet 0.1478 0.3028 0.2697 0.3199 0.3066 0.3453 0.3167 0.3340 0.3274 0.0854 0.0545 0.3531 0.3270
XGBoost 0.4494 0.4494 0.4494 0.4494 0.4494 0.4494 0.4494 0.4494 0.4494 0.4494 0.4494 0.4494 0.4494
TABLE XV: $R^{2}$ Score by dataset, model, and scaling method.
TABLE XVI: MSE by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
Abalone Ada 8.4579 7.9106 7.8773 7.6131 8.6210 8.1869 8.8383 7.8177 8.7869 8.1054 8.4388 8.0102 8.0102
CART 8.7504 8.6683 8.6547 8.5702 8.6762 8.6316 8.5032 8.7161 8.4657 8.7169 8.7600 8.4426 8.4426
CatBoost 4.8477 4.8482 4.8485 4.8478 4.8486 4.8488 4.8492 4.8470 4.8501 4.8480 4.8477 4.8505 4.8506
KNN 4.9107 5.0541 5.1229 5.4200 5.0202 5.6801 4.9107 5.5324 5.3144 5.9916 5.4200 5.3829 5.3829
LGBM 4.8137 4.7948 4.8137 4.8170 4.8175 4.8175 4.8175 4.8846 4.7960 4.8137 4.8158 4.8158 4.8146
LinearRegression 4.9249 4.9249 4.9249 4.9249 4.9249 4.9249 4.9249 4.9249 5.1136 4.9249 4.9242 4.8700 4.8700
MLP 4.8285 4.7973 4.8087 4.4900 4.4885 4.6276 4.5099 4.4356 4.6370 4.7730 9.9246 4.7351 4.5088
RF 4.8295 4.8261 4.8244 4.8392 4.8299 4.8257 4.8268 4.8328 4.8436 4.8421 4.8342 4.8376 4.8376
SVR 4.7801 4.7898 4.8158 4.6502 4.5642 4.9178 4.6401 4.6734 4.7267 6.4835 4.6502 4.6337 4.6337
TabNet 6.9789 6.4898 8.0072 4.9438 4.8501 4.5143 6.3516 4.6369 5.9460 10.7584 10.6238 6.6442 5.0601
XGBoost 5.5381 5.5381 5.5381 5.5381 5.5381 5.5381 5.5381 5.5381 5.5381 5.5381 5.5381 5.5381 5.5381
Air Quality Ada 1.3054 1.3054 2.8180 2.8180 2.8150 2.7004 1.3054 1.3054 2.8118 1.8197 2.8151 2.7332 2.7332
CART 0.0264 0.0264 0.0264 0.0264 0.0264 0.0751 0.0264 0.0264 0.0264 0.0264 0.0265 0.0265 0.0265
CatBoost 0.1474 0.1475 0.1473 0.1473 0.1473 0.1768 0.1473 0.1473 0.1474 0.1473 0.1392 0.1478 0.1478
KNN 0.9109 1.1487 1.2492 1.0231 0.9524 0.6838 0.9109 2.3434 2.4706 1.2911 1.0232 1.0926 1.0926
LGBM 0.1129 0.1144 0.1129 0.1298 0.1298 0.1332 0.1298 0.1291 0.1144 0.1129 0.1156 0.1156 0.1300
LinearRegression 1.4180 1.4180 1.4180 1.4180 1.4180 1.4180 1.4180 1.4180 217.6832 1.4180 1.4189 3.9983 3.9983
MLP 2.6090 1.6711 0.0278 0.0110 0.0910 0.0229 4.2237 0.1181 2.8912 1.4552 312.0822 0.5080 0.1000
RF 0.0131 0.0133 0.0131 0.0131 0.0131 0.0126 0.0131 0.0131 0.0104 0.0132 0.0137 0.0130 0.0130
SVR 5.7798 35.8845 65.6306 125.8177 97.2254 14.6127 55.9688 139.7765 387.6257 79.1870 125.8348 121.0407 121.0407
TabNet 0.5921 2.7576 0.8953 3.7794 0.4268 3.2731 1.2790 0.3880 8.9527 22.4788 1661.0442 0.3756 1.6444
XGBoost 0.1147 0.1147 0.1147 0.1147 0.1147 0.1079 0.1147 0.1147 0.1147 0.1147 0.1364 0.1147 0.1147
Appliances Energy Prediction Ada 37427.0812 18498.3970 49715.5179 45763.4907 51370.2541 48177.6281 25307.3469 25307.3469 28453.4699 18498.3970 25305.0426 18498.3970 18498.3970
CART 8903.9689 8913.8490 8900.0338 8882.6043 8883.1110 8894.7137 8907.5494 8884.2763 8716.1797 8854.3151 8862.6921 8849.8564 8849.8564
CatBoost 5636.6122 5594.4777 5636.8339 5594.4722 5594.4462 5594.5548 5594.4195 5594.4042 5575.4147 5636.6466 5596.2357 5594.2071 5594.0726
KNN 8570.5671 7340.9715 8192.0770 6924.2331 7629.3687 5439.7190 8570.5671 7285.3653 6977.3282 8687.5352 6924.2331 6922.9272 6922.9272
LGBM 5854.3508 5897.1553 5854.3508 5837.8269 5837.8269 5837.8269 5837.8269 5983.4385 5887.6446 5854.3508 5819.6456 5819.6456 5837.9645
LinearRegression 8579.6535 8579.6535 8579.6535 8579.6535 8579.6535 8579.6535 8579.6535 8579.6535 8769.7261 8579.6535 8579.5634 8609.0530 8609.0530
MLP 8656.1894 8203.8308 8673.7742 7063.7610 7116.0726 7046.2702 7194.4738 7243.1268 7992.0693 8715.8666 10299.4889 8027.0389 7277.7891
RF 5025.6175 5025.6061 5028.0530 5025.4432 5018.0357 5023.1468 5020.2272 5022.6680 5033.6409 5015.6727 5027.8705 5021.9574 5021.2540
SVR 11390.9162 10335.1997 10585.9003 10144.3973 10259.9281 10836.6115 10459.6290 10401.7762 10198.0740 10724.4861 10144.4020 10189.3769 10189.3769
TabNet 7156.6395 7258.0613 7029.5273 7389.1339 7180.2959 6868.7284 7160.9294 7443.5996 7215.4305 7378.4224 6946.2734 7239.4565 7010.9228
XGBoost 5369.4787 5369.4787 5369.4787 5369.4787 5369.4787 5369.4787 5369.4787 5369.4787 5314.8232 5369.4787 5369.4846 5369.4787 5369.4787
Concrete Compressive Strength Ada 60.5643 60.5890 61.7613 60.0204 59.9916 59.8600 58.7378 59.8600 59.6016 59.8600 59.1379 57.5319 57.5319
CART 47.2904 47.2904 47.2990 47.2990 47.2904 47.2904 47.3263 47.3263 49.0908 47.2904 47.3263 47.3510 47.3510
CatBoost 17.9329 17.9310 17.9327 17.9331 17.9330 17.9158 17.9344 17.9199 17.7440 17.9171 17.9348 17.9007 17.9007
KNN 87.4024 95.4699 91.1487 88.9093 79.6278 139.2635 87.3197 69.1034 54.8975 97.9166 88.9093 77.9513 77.9513
LGBM 20.8520 20.8520 20.8520 21.1791 21.1791 21.1791 21.1791 20.9335 20.7375 20.8520 20.7807 20.7807 21.1283
LinearRegression 109.7508 109.7508 109.7508 109.7508 109.7508 109.7508 109.7508 109.7508 52.5905 109.7508 109.7231 81.6210 81.6210
MLP 53.3124 56.2212 68.5215 34.5083 33.9975 45.4931 53.9233 36.1983 45.1127 116.8063 271.2155 69.5416 44.8916
RF 29.8643 29.9020 29.8945 30.0114 29.9181 29.8971 29.9510 29.9231 29.6500 29.8871 29.8991 29.8406 29.8406
SVR 209.4406 119.2487 124.6258 105.7187 111.0081 168.9885 123.2218 81.5215 72.9884 164.4260 105.7134 99.7138 99.7138
TabNet 2245230.0052 1498.7334 1496.2369 1448.4442 2424.1084 1312.0591 237345.3437 1477.8487 1507.0072 1494.0005 1536.9135 1516.1655 1478.5968
XGBoost 24.2314 24.2314 24.2314 24.2314 24.2314 24.2314 24.2314 24.2314 24.2314 24.2314 24.2314 24.2314 24.2314
Forest Fires Ada 11920.6614 11818.4033 11483.3932 8401.7532 8457.8291 11713.7349 11728.3039 11469.2245 8325.6368 11594.6689 11804.1047 8109.8741 8109.8741
CART 13278.1984 13284.7052 13284.7052 13284.7052 13252.0296 13278.1984 13284.7052 13284.7052 13076.4297 13073.6727 13284.7052 13068.9608 13068.9608
CatBoost 8305.6236 8305.5795 8305.5795 8305.5795 8305.8724 8305.5695 8306.0607 8305.5795 8306.2021 8305.5983 8305.5994 8305.8919 8305.8919
KNN 8049.5193 8450.3037 8331.5751 8314.0348 7683.6465 8380.0861 8049.5193 8232.2422 7874.0009 8394.1118 8314.0348 8278.8926 8278.8926
LGBM 8153.9240 8153.9240 8153.9240 8107.4363 8107.4363 8107.4363 8107.4363 8065.4234 8187.3561 8153.9240 8101.9144 8101.9144 8058.7846
LinearRegression 7925.6571 7925.6571 7925.6571 7925.6571 7925.6571 7925.6571 7925.6571 7925.6571 7884.4617 7925.6571 7925.6456 7909.6969 7909.6969
MLP 8010.9478 7903.1906 7912.5614 7947.6005 8169.4083 7963.7329 8339.4658 7861.6344 7851.0386 7908.2283 7967.3473 7886.8273 7843.6319
RF 8801.7409 8820.9222 8833.7029 8827.5833 8827.4812 8749.3543 8828.6643 8799.6126 8757.0734 8835.0585 8801.6836 8801.2509 8801.2509
SVR 8162.5768 8153.9647 8154.0176 8152.2006 8161.5094 8155.5670 8168.8605 8152.5913 8155.7694 8158.6271 8152.2076 8150.3462 8150.3462
TabNet 17388.5255 8175.7646 8167.9740 8161.3423 8304.3410 8197.3670 18285.8949 8166.9890 8171.4319 8177.1570 8167.8719 8164.7159 8174.7450
XGBoost 11685.4576 11685.4576 11685.4576 11685.4576 11685.4576 11685.4576 11685.4576 11685.4576 11685.4576 11685.4576 11685.4576 11685.4576 11685.4576
Real Estate Valuation Ada 57.2065 52.7882 52.7882 53.2506 52.7882 54.8695 54.7434 56.1439 52.8036 50.5103 55.5551 53.0226 53.0226
CART 73.4509 73.1951 73.1951 73.1951 73.4537 72.9267 73.4537 73.4509 74.1349 66.7002 73.1951 74.1321 74.1321
CatBoost 44.7412 44.7336 44.7336 44.7413 44.7257 44.7387 44.7055 44.7387 45.1346 44.7321 44.7055 44.7879 44.7879
KNN 63.0027 63.7030 63.0182 64.3252 58.7216 88.8102 63.0027 61.0770 51.4310 72.8293 64.3252 62.6561 62.6561
LGBM 50.1547 50.1547 50.1547 48.1590 48.1590 48.1590 48.1590 48.9211 50.1370 50.1547 50.1674 50.1674 48.1683
LinearRegression 73.5684 73.5684 73.5684 73.5684 73.5684 73.5684 73.5684 73.5684 53.4317 73.5684 73.5625 66.0181 66.0181
MLP 63.5585 70.8757 73.4388 59.6005 56.1629 960.9785 118.4726 51.9328 60.4070 80.5388 168.5185 73.8182 57.9370
RF 42.7505 42.6567 42.6657 42.7479 42.8437 42.6366 42.7353 42.7405 41.5067 42.5372 42.6543 42.9702 42.9702
SVR 85.3424 73.4157 78.1453 70.4354 75.5672 98.7492 80.2402 69.0369 61.6845 80.2676 70.4315 67.7575 67.7575
TabNet 3934014.4970 1621.3258 1643.8788 1621.5964 1583.2938 355272072.3309 216631.9725 1578.2590 1610.2450 1546.2204 1595.7477 1616.1390 1552.9722
XGBoost 49.9527 49.9527 49.9527 49.9527 49.9527 49.9527 49.9527 49.9527 49.9527 49.9527 49.9527 49.9527 49.9527
Wine Quality Ada 0.5040 0.5004 0.5098 0.5095 0.5131 0.5109 0.5112 0.4992 0.5139 0.5092 0.4999 0.4999 0.4999
CART 0.7072 0.7128 0.7097 0.7087 0.7123 0.7026 0.7036 0.7062 0.7210 0.7123 0.7092 0.7149 0.7149
CatBoost 0.3862 0.3862 0.3862 0.3862 0.3862 0.3861 0.3862 0.3862 0.3862 0.3862 0.3862 0.3860 0.3860
KNN 0.6406 0.5141 0.4901 0.4761 0.5322 0.6217 0.6406 0.4937 0.4929 0.5011 0.4762 0.4818 0.4818
LGBM 0.3971 0.3971 0.3971 0.3956 0.3956 0.3956 0.3956 0.3939 0.3972 0.3971 0.3973 0.3973 0.3956
LinearRegression 0.5326 0.5326 0.5326 0.5326 0.5326 0.5326 0.5326 0.5326 0.5310 0.5326 0.5325 0.5245 0.5245
MLP 0.5472 0.4775 0.4930 0.4471 0.4587 0.5459 0.5089 0.4466 0.4704 0.5160 0.7315 0.4856 0.4477
RF 0.3659 0.3658 0.3655 0.3661 0.3653 0.3658 0.3655 0.3654 0.3658 0.3662 0.3664 0.3654 0.3654
SVR 0.6149 0.4698 0.4973 0.4493 0.4874 0.5753 0.5629 0.4525 0.4505 0.5014 0.4493 0.4446 0.4446
TabNet 0.6218 0.5087 0.5329 0.4963 0.5059 0.4777 0.4986 0.4859 0.4908 0.6674 0.6899 0.4720 0.4910
XGBoost 0.4017 0.4017 0.4017 0.4017 0.4017 0.4017 0.4017 0.4017 0.4017 0.4017 0.4017 0.4017 0.4017
TABLE XVI: MSE by dataset, model, and scaling method.
TABLE XVII: MAE by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
Abalone Ada 2.5042 2.4037 2.3991 2.3437 2.5275 2.4467 2.5691 2.3881 2.5624 2.4383 2.4968 2.4173 2.4173
CART 2.0805 2.0702 2.0710 2.0582 2.0766 2.0702 2.0582 2.0718 2.0542 2.0789 2.0821 2.0502 2.0502
CatBoost 1.5612 1.5614 1.5612 1.5612 1.5617 1.5611 1.5616 1.5603 1.5620 1.5611 1.5609 1.5618 1.5618
KNN 1.5673 1.5938 1.6102 1.6555 1.5943 1.6962 1.5673 1.6654 1.6140 1.7400 1.6555 1.6416 1.6416
LGBM 1.5476 1.5486 1.5476 1.5500 1.5502 1.5502 1.5502 1.5632 1.5489 1.5476 1.5484 1.5484 1.5494
LinearRegression 1.6187 1.6187 1.6187 1.6187 1.6187 1.6187 1.6187 1.6187 1.6734 1.6187 1.6186 1.6132 1.6132
MLP 1.6159 1.5740 1.5748 1.5318 1.5059 1.5392 1.5258 1.4958 1.5429 1.5844 2.3339 1.5816 1.5126
RF 1.5590 1.5595 1.5584 1.5617 1.5596 1.5601 1.5593 1.5619 1.5613 1.5615 1.5622 1.5613 1.5613
SVR 1.5048 1.5086 1.5111 1.4964 1.4830 1.5272 1.4856 1.4990 1.4972 1.6810 1.4963 1.4900 1.4900
TabNet 2.0912 1.9773 2.3117 1.5785 1.6197 1.4921 1.9571 1.4957 1.8579 2.5043 2.5471 2.0374 1.7154
XGBoost 1.6623 1.6623 1.6623 1.6623 1.6623 1.6623 1.6623 1.6623 1.6623 1.6623 1.6623 1.6623 1.6623
Air Quality Ada 0.8698 0.8698 1.3721 1.3720 1.3685 1.3244 0.8698 0.8698 1.3417 1.0721 1.3685 1.3383 1.3383
CART 0.0203 0.0203 0.0202 0.0202 0.0202 0.0235 0.0203 0.0202 0.0200 0.0202 0.0203 0.0203 0.0203
CatBoost 0.1711 0.1712 0.1710 0.1711 0.1711 0.1863 0.1711 0.1711 0.1712 0.1711 0.1710 0.1714 0.1714
KNN 0.5506 0.6348 0.6755 0.5862 0.5577 0.4738 0.5506 0.8379 0.7639 0.6980 0.5864 0.5759 0.5759
LGBM 0.0677 0.0669 0.0677 0.0728 0.0728 0.0731 0.0728 0.0731 0.0669 0.0677 0.0678 0.0678 0.0730
LinearRegression 0.8282 0.8282 0.8282 0.8282 0.8282 0.8282 0.8282 0.8282 10.9540 0.8282 0.8284 1.2383 1.2383
MLP 1.2505 0.8289 0.0909 0.0698 0.2038 0.0988 1.2563 0.1910 0.8323 0.8479 13.5283 0.3256 0.1266
RF 0.0167 0.0169 0.0167 0.0166 0.0167 0.0169 0.0167 0.0167 0.0159 0.0168 0.0168 0.0166 0.0166
SVR 0.8970 1.2103 1.4802 1.9293 1.7292 0.6919 1.3269 5.4424 4.7159 1.4842 1.9294 1.9787 1.9787
TabNet 0.4924 0.9669 0.6923 0.9567 0.4297 0.5142 0.6516 0.4076 1.1264 3.1115 12.9536 0.3352 0.7997
XGBoost 0.0769 0.0769 0.0769 0.0769 0.0769 0.0764 0.0769 0.0769 0.0769 0.0769 0.0790 0.0769 0.0769
Appliances Energy Prediction Ada 183.0879 124.1897 212.6930 203.8061 216.6901 209.1780 148.9915 148.9915 157.4692 124.1897 148.9851 124.1897 124.1897
CART 40.5540 40.6029 40.5692 40.4712 40.5455 40.5404 40.5573 40.5269 40.3091 40.4087 40.3952 40.4256 40.4256
CatBoost 38.3286 38.2656 38.3309 38.2655 38.2655 38.2659 38.2642 38.2649 38.1813 38.3277 38.2543 38.2625 38.2620
KNN 47.7696 42.0912 45.5977 39.8189 43.2660 34.0929 47.7696 41.5828 41.1120 48.6708 39.8189 40.4084 40.4084
LGBM 39.0633 39.0817 39.0633 39.0969 39.0969 39.0969 39.0969 39.3198 39.2256 39.0633 39.1020 39.1020 39.0960
LinearRegression 52.9227 52.9227 52.9227 52.9227 52.9227 52.9227 52.9227 52.9227 53.9764 52.9227 52.9226 53.2947 53.2947
MLP 53.8080 51.1479 54.1215 47.7318 47.5991 47.7552 49.2947 47.7236 51.0640 53.4270 60.5754 51.3053 48.1224
RF 34.2701 34.2698 34.2923 34.2876 34.2666 34.2645 34.2602 34.2838 34.3702 34.2628 34.2888 34.3111 34.3082
SVR 48.9182 44.1754 45.5442 43.4164 44.1483 46.1472 45.1832 45.1344 43.8597 46.2897 43.4165 43.6265 43.6265
TabNet 43.6365 44.2129 43.0738 44.9472 43.2715 42.8673 43.5117 45.0579 43.5039 44.3062 42.3802 44.3074 44.7821
XGBoost 37.3973 37.3973 37.3973 37.3973 37.3973 37.3973 37.3973 37.3973 37.3315 37.3973 37.3974 37.3973 37.3973
Concrete Compressive Strength Ada 6.4633 6.4051 6.5035 6.3843 6.4167 6.3515 6.3480 6.3515 6.3690 6.3515 6.3553 6.2855 6.2855
CART 4.4604 4.4604 4.4623 4.4623 4.4604 4.4604 4.4647 4.4647 4.5501 4.4604 4.4647 4.4681 4.4681
CatBoost 2.7582 2.7570 2.7581 2.7579 2.7586 2.7574 2.7588 2.7559 2.7539 2.7575 2.7589 2.7529 2.7529
KNN 7.2301 7.6315 7.0405 7.3319 7.0126 9.5001 7.2247 6.4237 5.4080 7.3897 7.3319 6.8437 6.8437
LGBM 3.0480 3.0480 3.0480 3.0473 3.0473 3.0473 3.0473 3.0197 3.0303 3.0480 3.0353 3.0353 3.0364
LinearRegression 8.2986 8.2986 8.2986 8.2986 8.2986 8.2986 8.2986 8.2986 5.7572 8.2986 8.2973 7.0897 7.0897
MLP 5.9685 5.7620 6.4045 4.4758 4.2648 5.2214 5.3184 4.5392 5.3287 8.6028 13.3172 6.5977 5.1541
RF 3.7512 3.7531 3.7503 3.7608 3.7547 3.7539 3.7572 3.7550 3.7495 3.7511 3.7528 3.7430 3.7430
SVR 11.6674 8.6790 8.9978 8.1314 8.2733 10.7062 8.6913 7.0217 6.5147 10.4403 8.1311 7.8360 7.8360
TabNet 1149.8143 35.0352 35.0250 34.2232 39.0829 31.0733 268.8924 34.6500 35.1275 34.9598 35.5915 35.2705 34.6666
XGBoost 3.1739 3.1739 3.1739 3.1739 3.1739 3.1739 3.1739 3.1739 3.1739 3.1739 3.1739 3.1739 3.1739
Forest Fires Ada 34.7652 31.7109 36.7437 29.3869 29.5476 36.1184 43.9637 33.7898 34.3625 35.6878 32.0355 28.5811 28.5811
CART 32.0911 32.2953 32.2953 32.2953 31.8315 32.0911 32.2953 32.2953 31.0349 30.9868 32.2953 30.9321 30.9321
CatBoost 21.9476 21.9470 21.9470 21.9470 21.9597 21.9468 21.9593 21.9470 21.9818 21.9501 21.9498 21.9440 21.9440
KNN 21.5065 20.9572 20.7837 21.0229 21.7229 20.3066 21.5065 19.7818 18.7445 20.2599 21.0229 20.6635 20.6635
LGBM 24.2821 24.2821 24.2821 24.1922 24.1922 24.1922 24.1922 23.9371 24.5966 24.2821 23.7689 23.7689 23.7599
LinearRegression 20.7980 20.7980 20.7980 20.7980 20.7980 20.7980 20.7980 20.7980 20.8690 20.7980 20.7983 20.9673 20.9673
MLP 21.2084 20.5521 20.6192 24.5951 24.6872 24.3192 24.9473 23.9833 20.7647 20.6168 20.1942 20.7236 23.1578
RF 24.2812 24.2711 24.3738 24.3496 24.3699 23.8711 24.3590 24.3443 24.0253 24.3326 24.4005 24.3836 24.3836
SVR 14.9474 14.9723 14.9755 14.9655 14.9625 14.9637 14.9233 14.9590 15.0261 15.0110 14.9655 14.9942 14.9942
TabNet 94.6647 14.9556 14.9436 15.1497 20.5845 17.4966 61.7420 15.0331 14.9516 14.9664 14.9649 14.9054 14.9939
XGBoost 29.0770 29.0770 29.0770 29.0770 29.0770 29.0770 29.0770 29.0770 29.0770 29.0770 29.0770 29.0770 29.0770
Real Estate Valuation Ada 5.6034 5.3289 5.3289 5.4391 5.3289 5.5175 5.5500 5.5659 5.3031 5.1177 5.6168 5.4525 5.4525
CART 5.6076 5.5892 5.5892 5.5892 5.6116 5.5868 5.6116 5.6076 5.6700 5.3548 5.5892 5.6660 5.6660
CatBoost 4.4727 4.4703 4.4703 4.4735 4.4681 4.4711 4.4642 4.4711 4.5069 4.4667 4.4642 4.4764 4.4764
KNN 5.4626 5.6760 5.4435 5.6979 5.3102 6.7592 5.4626 5.5819 5.1686 5.8458 5.6979 5.5866 5.5866
LGBM 4.8106 4.8106 4.8106 4.7284 4.7284 4.7284 4.7284 4.8254 4.8065 4.8106 4.8187 4.8187 4.7326
LinearRegression 6.1848 6.1848 6.1848 6.1848 6.1848 6.1848 6.1848 6.1848 4.9362 6.1848 6.1844 5.7043 5.7043
MLP 5.3601 6.1315 6.1126 5.4400 5.0396 21.0560 8.8945 5.0138 5.4359 6.5674 10.7036 6.3407 5.3635
RF 4.3971 4.3919 4.3951 4.4002 4.4044 4.3726 4.4000 4.3916 4.3922 4.3867 4.3892 4.4019 4.4019
SVR 6.8824 6.1177 6.2845 5.9805 6.1307 7.6388 6.5489 5.8563 5.5097 6.3822 5.9803 5.9063 5.9063
TabNet 1625.5671 38.0836 38.2156 37.9038 37.8126 14245.0982 271.3981 37.6000 37.9313 37.1465 37.7939 38.0048 37.1589
XGBoost 4.7311 4.7311 4.7311 4.7311 4.7311 4.7311 4.7311 4.7311 4.7311 4.7311 4.7311 4.7311 4.7311
Wine Quality Ada 0.5608 0.5751 0.5633 0.5607 0.5719 0.5696 0.5622 0.5606 0.5686 0.5634 0.5612 0.5568 0.5568
CART 0.4928 0.4944 0.4933 0.4923 0.4959 0.4913 0.4913 0.4928 0.4985 0.4949 0.4928 0.4954 0.4954
CatBoost 0.4796 0.4795 0.4796 0.4796 0.4796 0.4796 0.4796 0.4796 0.4797 0.4796 0.4796 0.4796 0.4796
KNN 0.6243 0.5406 0.5231 0.5259 0.5558 0.5967 0.6243 0.5349 0.5275 0.5350 0.5261 0.5280 0.5280
LGBM 0.4832 0.4832 0.4832 0.4829 0.4829 0.4829 0.4829 0.4833 0.4832 0.4832 0.4833 0.4833 0.4829
LinearRegression 0.5639 0.5639 0.5639 0.5639 0.5639 0.5639 0.5639 0.5639 0.5627 0.5639 0.5639 0.5617 0.5617
MLP 0.5692 0.5436 0.5499 0.5187 0.5299 0.5742 0.5563 0.5196 0.5376 0.5598 0.6657 0.5471 0.5261
RF 0.4366 0.4367 0.4368 0.4370 0.4364 0.4365 0.4364 0.4362 0.4371 0.4370 0.4369 0.4365 0.4365
SVR 0.6076 0.5254 0.5453 0.5111 0.5454 0.5921 0.5893 0.5142 0.5080 0.5480 0.5111 0.5065 0.5065
TabNet 0.6061 0.5582 0.5779 0.5453 0.5494 0.5324 0.5574 0.5428 0.5483 0.6509 0.6585 0.5386 0.5515
XGBoost 0.4639 0.4639 0.4639 0.4639 0.4639 0.4639 0.4639 0.4639 0.4639 0.4639 0.4639 0.4639 0.4639
TABLE XVII: MAE by dataset, model, and scaling method.
TABLE XVIII: Time to Train (s) by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
Abalone Ada 0.0971 0.1051 0.0774 0.0758 0.0961 0.0966 0.0967 0.0810 0.0971 0.0979 0.0973 0.1004 0.0978
CART 0.0095 0.0093 0.0093 0.0093 0.0094 0.0094 0.0093 0.0093 0.0093 0.0095 0.0094 0.0094 0.0094
CatBoost 0.6395 0.6556 0.6344 0.6431 0.6465 0.6445 0.6430 0.6503 0.6530 0.6478 0.6459 0.6400 0.6432
KNN 0.0010 0.0011 0.0010 0.0010 0.0010 0.0010 0.0010 0.0010 0.0010 0.0010 0.0011 0.0011 0.0011
LGBM 0.0286 0.0290 0.0288 0.0301 0.0303 0.0304 0.0304 0.0307 0.0293 0.0297 0.0292 0.0283 0.0303
LinearRegression 0.0003 0.0091 0.0004 0.0003 0.0004 0.0004 0.0003 0.0003 0.0003 0.0004 0.0004 0.0004 0.0004
MLP 0.8414 1.1040 1.1404 0.8932 1.1691 0.7039 1.5363 1.2343 1.8627 1.7507 0.3577 1.4888 1.5986
RF 0.5945 0.6241 0.6035 0.6003 0.5983 0.6006 0.5963 0.5977 0.6015 0.5973 0.6022 0.5991 0.5997
SVR 0.1382 0.1348 0.1383 0.1392 0.1382 0.1383 0.1397 0.1440 0.1391 0.1367 0.1369 0.1367 0.1372
TabNet 2.8896 2.9717 2.8425 2.7805 2.7096 2.6584 2.7499 2.8266 2.6811 2.9476 2.7365 2.8925 2.7076
XGBoost 0.0287 0.0291 0.0303 0.0290 0.0296 0.0300 0.0289 0.0305 0.0284 0.0284 0.0286 0.0286 0.0285
Air Quality Ada 0.2988 0.3068 0.1755 0.1752 0.2095 0.1829 0.3000 0.2992 0.1339 0.3224 0.2055 0.1709 0.1701
CART 0.0297 0.0293 0.0296 0.0293 0.0294 0.0294 0.0292 0.0295 0.0292 0.0295 0.0289 0.0292 0.0296
CatBoost 0.9429 0.9439 0.9445 0.9462 0.9445 0.9468 0.9331 0.9437 0.9376 0.9450 0.9409 0.9413 0.9282
KNN 0.0032 0.0033 0.0032 0.0031 0.0032 0.0032 0.0033 0.0031 0.0032 0.0035 0.0032 0.0034 0.0034
LGBM 0.0391 0.0384 0.0395 0.0392 0.0388 0.0390 0.0393 0.0396 0.0391 0.0405 0.0390 0.0377 0.0391
LinearRegression 0.0006 0.0008 0.0006 0.0005 0.0006 0.0006 0.0006 0.0006 0.0006 0.0006 0.0006 0.0006 0.0006
MLP 0.5701 1.8147 3.3321 2.1864 1.9174 2.4677 1.5022 1.8880 2.8593 4.4512 6.4014 5.9718 4.9511
RF 1.8482 1.8810 1.8664 1.8547 1.8550 1.8489 1.8500 1.8659 1.9096 1.8457 1.8309 1.8409 1.8465
SVR 0.6315 0.5922 0.5742 0.4535 0.4049 0.3582 0.4079 0.7643 0.5973 0.5838 0.4433 0.4812 0.4799
TabNet 8.0233 8.2011 8.0828 8.2231 8.0882 8.0729 8.3433 7.9980 7.9860 8.2220 8.2484 8.3199 8.1277
XGBoost 0.0419 0.0422 0.0438 0.0431 0.0431 0.0415 0.0421 0.0450 0.0417 0.0419 0.0412 0.0418 0.0418
Appliances Energy Prediction Ada 1.1052 0.5338 1.6459 1.4369 1.6431 1.6525 0.6962 0.6970 0.7305 0.5321 0.7032 0.5343 0.5393
CART 0.2910 0.2860 0.2859 0.2859 0.2854 0.2859 0.2864 0.2866 0.2854 0.2863 0.2861 0.2857 0.2876
CatBoost 1.3813 1.3763 1.3719 1.3633 1.3738 1.3676 1.3759 1.3649 1.3771 1.3725 1.3655 1.3775 1.3878
KNN 0.0004 0.0004 0.0004 0.0004 0.0005 0.0004 0.0004 0.0004 0.0004 0.0005 0.0004 0.0004 0.0004
LGBM 0.0522 0.0508 0.0532 0.0552 0.0551 0.0553 0.0553 0.0545 0.0520 0.0533 0.0527 0.0511 0.0580
LinearRegression 0.0013 0.0020 0.0014 0.0015 0.0015 0.0016 0.0016 0.0014 0.0014 0.0017 0.0016 0.0016 0.0016
MLP 3.3408 14.3606 12.6484 17.8381 16.0423 10.0974 12.9527 16.1881 14.7886 15.5263 0.3227 15.9089 16.6124
RF 18.2580 18.4922 18.3987 18.3023 18.3270 18.3389 18.2072 18.2253 18.5976 18.2453 18.2356 18.5935 18.3139
SVR 3.6868 3.6235 3.6063 3.6767 3.5467 3.7634 3.5353 3.7207 3.7124 3.5425 3.5711 3.5809 3.5765
TabNet 17.4699 17.6083 17.7432 17.2341 17.3906 17.3073 17.9814 17.3632 17.5384 17.8526 17.6793 17.8897 17.6663
XGBoost 0.0772 0.0779 0.0818 0.0820 0.0787 0.0774 0.0773 0.0821 0.0764 0.0772 0.0773 0.0774 0.0769
Concrete Compressive Strength Ada 0.0399 0.0397 0.0397 0.0400 0.0395 0.0396 0.0396 0.0398 0.0398 0.0397 0.0414 0.0403 0.0402
CART 0.0020 0.0020 0.0020 0.0027 0.0020 0.0020 0.0020 0.0021 0.0020 0.0020 0.0020 0.0020 0.0021
CatBoost 0.4633 0.4593 0.4706 0.4671 0.4660 0.4733 0.4634 0.4705 0.4645 0.4675 0.4638 0.4664 0.4646
KNN 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004
LGBM 0.0221 0.0227 0.0232 0.0231 0.0231 0.0264 0.0236 0.0237 0.0231 0.0230 0.0230 0.0222 0.0245
LinearRegression 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0004 0.0004
MLP 0.1335 0.7523 0.7281 0.8502 0.8581 0.8483 0.6685 0.8323 0.7186 0.7916 0.1296 0.7726 0.8178
RF 0.1397 0.1399 0.1396 0.1390 0.1389 0.1388 0.1387 0.1438 0.1392 0.1399 0.1385 0.1390 0.1388
SVR 0.0086 0.0083 0.0093 0.0087 0.0084 0.0098 0.0086 0.0091 0.0087 0.0083 0.0084 0.0084 0.0083
TabNet 0.0348 0.0342 0.0356 0.0348 0.0346 0.0355 0.0357 0.0343 0.0365 0.0357 0.0356 0.0361 0.0345
XGBoost 0.0253 0.0242 0.0261 0.0265 0.0249 0.0243 0.0243 0.0257 0.0237 0.0241 0.0239 0.0240 0.0237
Forest Fires Ada 0.0200 0.0117 0.0193 0.0121 0.0141 0.0302 0.0302 0.0153 0.0218 0.0143 0.0148 0.0139 0.0139
CART 0.0014 0.0014 0.0016 0.0014 0.0015 0.0014 0.0014 0.0014 0.0015 0.0015 0.0015 0.0014 0.0015
CatBoost 0.3892 0.3886 0.3922 0.3870 0.3903 0.3968 0.3836 0.3906 0.3869 0.3945 0.3834 0.3869 0.3951
KNN 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
LGBM 0.0115 0.0120 0.0121 0.0119 0.0119 0.0127 0.0117 0.0118 0.0122 0.0121 0.0121 0.0119 0.0116
LinearRegression 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
MLP 0.2457 0.3855 0.3816 0.4316 0.4790 0.4430 0.4089 0.4330 0.3801 0.3860 0.0644 0.3873 0.4248
RF 0.0997 0.1010 0.1006 0.1002 0.1006 0.1001 0.1002 0.1063 0.1001 0.1006 0.1007 0.1000 0.1005
SVR 0.0025 0.0026 0.0028 0.0027 0.0025 0.0026 0.0025 0.0025 0.0025 0.0024 0.0026 0.0026 0.0026
TabNet 0.0325 0.0303 0.0320 0.0361 0.0306 0.0308 0.0325 0.0308 0.0328 0.0325 0.0323 0.0340 0.0306
XGBoost 0.0215 0.0207 0.0205 0.0223 0.0206 0.0202 0.0202 0.0214 0.0197 0.0199 0.0198 0.0199 0.0198
Real Estate Valuation Ada 0.0268 0.0269 0.0267 0.0267 0.0267 0.0268 0.0269 0.0269 0.0269 0.0269 0.0282 0.0276 0.0270
CART 0.0008 0.0008 0.0009 0.0008 0.0009 0.0008 0.0009 0.0008 0.0008 0.0009 0.0009 0.0009 0.0009
CatBoost 0.3852 0.3966 0.3833 0.3865 0.3852 0.3937 0.3925 0.3936 0.3962 0.3872 0.3816 0.3950 0.3953
KNN 0.0002 0.0002 0.0002 0.0002 0.0003 0.0002 0.0002 0.0002 0.0002 0.0003 0.0002 0.0003 0.0002
LGBM 0.0090 0.0095 0.0094 0.0094 0.0097 0.0100 0.0094 0.0098 0.0095 0.0094 0.0094 0.0091 0.0096
LinearRegression 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
MLP 0.1022 0.3264 0.3386 0.3599 0.3951 0.1556 0.1754 0.3778 0.3339 0.3329 0.1273 0.3417 0.3463
RF 0.0664 0.0672 0.0666 0.0663 0.0666 0.0660 0.0665 0.0689 0.0664 0.0660 0.0667 0.0663 0.0665
SVR 0.0019 0.0017 0.0018 0.0023 0.0017 0.0017 0.0018 0.0017 0.0017 0.0017 0.0017 0.0017 0.0016
TabNet 0.0306 0.0295 0.0319 0.0342 0.0296 0.0295 0.0312 0.0311 0.0327 0.0328 0.0314 0.0328 0.0318
XGBoost 0.0205 0.0206 0.0193 0.0205 0.0200 0.0195 0.0193 0.0205 0.0190 0.0190 0.0190 0.0193 0.0190
Wine Quality Ada 0.1357 0.1484 0.1660 0.1705 0.1126 0.1520 0.1815 0.1810 0.1824 0.1809 0.1299 0.1360 0.1364
CART 0.0200 0.0199 0.0198 0.0198 0.0198 0.0198 0.0199 0.0199 0.0198 0.0200 0.0198 0.0198 0.0199
CatBoost 0.7668 0.7568 0.7679 0.7628 0.7745 0.7653 0.7652 0.7749 0.7634 0.7586 0.7743 0.7668 0.7617
KNN 0.0022 0.0022 0.0022 0.0023 0.0022 0.0021 0.0021 0.0022 0.0023 0.0022 0.0023 0.0024 0.0023
LGBM 0.0314 0.0317 0.0313 0.0322 0.0323 0.0324 0.0321 0.0334 0.0324 0.0325 0.0313 0.0310 0.0328
LinearRegression 0.0004 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005
MLP 0.4713 1.7105 1.3455 2.3311 1.5885 0.8618 0.4947 1.6345 1.5528 1.3377 0.0908 1.3502 1.9695
RF 1.2844 1.3053 1.2891 1.2968 1.2922 1.2920 1.2855 1.3414 1.2833 1.2865 1.2846 1.2828 1.2808
SVR 0.3166 0.3176 0.3181 0.3239 0.3157 0.3177 0.3194 0.3213 0.3149 0.3054 0.3211 0.3169 0.3190
TabNet 5.5797 5.4072 5.3061 5.4588 5.3787 5.3858 5.6280 5.7121 5.6199 5.4181 5.5159 5.3852 5.5306
XGBoost 0.0350 0.0351 0.0334 0.0342 0.0353 0.0335 0.0336 0.0351 0.0328 0.0332 0.0331 0.0331 0.0332
TABLE XVIII: Time to Train (s) by dataset, model, and scaling method.
TABLE XIX: Time to Inference (s) by dataset, model, and scaling method.
Dataset Model NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
Abalone Ada 0.0025 0.0026 0.0019 0.0019 0.0025 0.0025 0.0025 0.0020 0.0025 0.0025 0.0025 0.0025 0.0025
CART 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0001 0.0002 0.0001 0.0002 0.0002 0.0002 0.0002
CatBoost 0.0007 0.0008 0.0010 0.0009 0.0009 0.0010 0.0010 0.0008 0.0007 0.0009 0.0008 0.0008 0.0007
KNN 0.0034 0.0045 0.0048 0.0042 0.0045 0.0036 0.0035 0.0041 0.0042 0.0029 0.0044 0.0043 0.0044
LGBM 0.0006 0.0008 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007 0.0007
LinearRegression 0.0000 0.0001 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
MLP 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
RF 0.0110 0.0107 0.0111 0.0110 0.0110 0.0110 0.0110 0.0109 0.0111 0.0109 0.0110 0.0111 0.0112
SVR 0.0652 0.0653 0.0652 0.0656 0.0647 0.0647 0.0654 0.0666 0.0647 0.0646 0.0642 0.0643 0.0643
TabNet 0.0094 0.0095 0.0092 0.0091 0.0092 0.0092 0.0094 0.0094 0.0092 0.0093 0.0092 0.0095 0.0092
XGBoost 0.0004 0.0005 0.0004 0.0004 0.0004 0.0004 0.0004 0.0005 0.0005 0.0004 0.0004 0.0004 0.0004
Air Quality Ada 0.0028 0.0028 0.0016 0.0016 0.0019 0.0017 0.0028 0.0028 0.0012 0.0029 0.0018 0.0015 0.0016
CART 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002
CatBoost 0.0011 0.0012 0.0010 0.0012 0.0011 0.0010 0.0010 0.0012 0.0013 0.0011 0.0010 0.0012 0.0011
KNN 0.0171 0.0322 0.0317 0.0338 0.0215 0.0190 0.0173 0.0381 0.0289 0.0238 0.0346 0.0282 0.0276
LGBM 0.0010 0.0010 0.0009 0.0009 0.0009 0.0010 0.0009 0.0009 0.0009 0.0009 0.0009 0.0009 0.0009
LinearRegression 0.0000 0.0001 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
MLP 0.0005 0.0007 0.0006 0.0006 0.0006 0.0006 0.0005 0.0006 0.0006 0.0006 0.0006 0.0007 0.0006
RF 0.0169 0.0167 0.0171 0.0170 0.0171 0.0169 0.0171 0.0171 0.0173 0.0170 0.0170 0.0171 0.0167
SVR 0.2986 0.2736 0.2676 0.1981 0.1717 0.1479 0.1690 0.3599 0.2825 0.2772 0.1929 0.2070 0.2069
TabNet 0.0167 0.0165 0.0165 0.0166 0.0166 0.0166 0.0167 0.0166 0.0164 0.0171 0.0165 0.0170 0.0167
XGBoost 0.0005 0.0006 0.0006 0.0007 0.0007 0.0006 0.0006 0.0007 0.0006 0.0005 0.0005 0.0006 0.0005
Appliances Energy Prediction Ada 0.0053 0.0024 0.0084 0.0072 0.0089 0.0083 0.0034 0.0033 0.0035 0.0024 0.0035 0.0024 0.0026
CART 0.0008 0.0008 0.0008 0.0008 0.0008 0.0007 0.0008 0.0008 0.0007 0.0008 0.0008 0.0008 0.0008
CatBoost 0.0017 0.0016 0.0015 0.0017 0.0017 0.0014 0.0013 0.0016 0.0015 0.0016 0.0014 0.0016 0.0017
KNN 0.0222 0.0214 0.0215 0.0225 0.0210 0.0204 0.0217 0.0216 0.0256 0.0214 0.0211 0.0213 0.0206
LGBM 0.0017 0.0017 0.0017 0.0017 0.0017 0.0017 0.0016 0.0017 0.0017 0.0017 0.0016 0.0016 0.0020
LinearRegression 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
MLP 0.0012 0.0013 0.0012 0.0012 0.0016 0.0014 0.0013 0.0012 0.0012 0.0014 0.0013 0.0013 0.0012
RF 0.0694 0.0686 0.0697 0.0700 0.0694 0.0702 0.0693 0.0692 0.0700 0.0692 0.0701 0.0701 0.0696
SVR 1.8656 1.7988 1.8341 1.7995 1.8016 1.7985 1.7991 1.8159 1.8055 1.7986 1.7933 1.7972 1.7930
TabNet 0.0347 0.0346 0.0345 0.0343 0.0347 0.0347 0.0350 0.0346 0.0349 0.0349 0.0350 0.0347 0.0350
XGBoost 0.0008 0.0009 0.0009 0.0010 0.0010 0.0009 0.0009 0.0010 0.0008 0.0008 0.0009 0.0010 0.0009
Concrete Compressive Strength Ada 0.0015 0.0015 0.0015 0.0015 0.0015 0.0015 0.0015 0.0015 0.0015 0.0015 0.0016 0.0015 0.0015
CART 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
CatBoost 0.0006 0.0006 0.0005 0.0005 0.0005 0.0005 0.0005 0.0006 0.0005 0.0006 0.0005 0.0005 0.0007
KNN 0.0009 0.0009 0.0010 0.0010 0.0010 0.0007 0.0009 0.0010 0.0010 0.0008 0.0011 0.0010 0.0009
LGBM 0.0004 0.0004 0.0004 0.0004 0.0004 0.0005 0.0004 0.0004 0.0004 0.0004 0.0004 0.0004 0.0005
LinearRegression 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
MLP 0.0001 0.0002 0.0002 0.0001 0.0002 0.0002 0.0001 0.0001 0.0001 0.0001 0.0001 0.0002 0.0001
RF 0.0033 0.0034 0.0034 0.0032 0.0034 0.0033 0.0033 0.0034 0.0034 0.0033 0.0032 0.0033 0.0033
SVR 0.0044 0.0040 0.0043 0.0041 0.0041 0.0041 0.0041 0.0041 0.0041 0.0041 0.0040 0.0040 0.0040
TabNet 0.0043 0.0041 0.0040 0.0043 0.0043 0.0044 0.0043 0.0041 0.0043 0.0043 0.0041 0.0045 0.0044
XGBoost 0.0003 0.0003 0.0004 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
Forest Fires Ada 0.0008 0.0005 0.0008 0.0005 0.0006 0.0013 0.0013 0.0007 0.0009 0.0006 0.0007 0.0006 0.0006
CART 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
CatBoost 0.0005 0.0004 0.0005 0.0004 0.0004 0.0005 0.0005 0.0004 0.0006 0.0004 0.0004 0.0005 0.0005
KNN 0.0003 0.0005 0.0006 0.0006 0.0004 0.0005 0.0004 0.0005 0.0006 0.0005 0.0007 0.0006 0.0006
LGBM 0.0003 0.0003 0.0004 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
LinearRegression 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
MLP 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
RF 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018 0.0018 0.0019
SVR 0.0011 0.0011 0.0012 0.0011 0.0011 0.0011 0.0011 0.0013 0.0012 0.0011 0.0011 0.0011 0.0011
TabNet 0.0040 0.0036 0.0037 0.0044 0.0037 0.0037 0.0039 0.0039 0.0037 0.0039 0.0036 0.0039 0.0049
XGBoost 0.0003 0.0004 0.0003 0.0004 0.0003 0.0003 0.0003 0.0004 0.0003 0.0003 0.0003 0.0003 0.0003
Real Estate Valuation Ada 0.0012 0.0013 0.0012 0.0013 0.0012 0.0013 0.0013 0.0013 0.0012 0.0013 0.0013 0.0013 0.0012
CART 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
CatBoost 0.0005 0.0004 0.0005 0.0005 0.0005 0.0004 0.0004 0.0005 0.0005 0.0004 0.0005 0.0004 0.0004
KNN 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0004 0.0006 0.0003
LGBM 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
LinearRegression 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
MLP 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
RF 0.0017 0.0017 0.0017 0.0017 0.0017 0.0017 0.0017 0.0017 0.0017 0.0017 0.0017 0.0017 0.0017
SVR 0.0007 0.0007 0.0009 0.0007 0.0007 0.0007 0.0007 0.0007 0.0008 0.0007 0.0007 0.0007 0.0007
TabNet 0.0033 0.0033 0.0034 0.0035 0.0034 0.0034 0.0035 0.0036 0.0034 0.0035 0.0035 0.0034 0.0041
XGBoost 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003 0.0003
Wine Quality Ada 0.0024 0.0026 0.0030 0.0031 0.0019 0.0027 0.0034 0.0034 0.0033 0.0034 0.0022 0.0025 0.0026
CART 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002 0.0002
CatBoost 0.0011 0.0012 0.0011 0.0009 0.0010 0.0009 0.0009 0.0009 0.0011 0.0010 0.0011 0.0009 0.0007
KNN 0.0048 0.0286 0.0257 0.0369 0.0086 0.0042 0.0047 0.0329 0.0321 0.0169 0.0385 0.0321 0.0310
LGBM 0.0008 0.0008 0.0008 0.0008 0.0008 0.0008 0.0008 0.0009 0.0008 0.0008 0.0008 0.0008 0.0008
LinearRegression 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000
MLP 0.0004 0.0004 0.0004 0.0004 0.0004 0.0005 0.0004 0.0005 0.0004 0.0005 0.0004 0.0004 0.0004
RF 0.0143 0.0148 0.0147 0.0145 0.0147 0.0145 0.0143 0.0151 0.0147 0.0147 0.0146 0.0146 0.0143
SVR 0.1501 0.1434 0.1527 0.1419 0.1478 0.1521 0.1517 0.1422 0.1396 0.1503 0.1413 0.1406 0.1406
TabNet 0.0116 0.0115 0.0115 0.0115 0.0115 0.0114 0.0120 0.0115 0.0118 0.0117 0.0120 0.0116 0.0117
XGBoost 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005 0.0005
TABLE XIX: Time to Inference (s) by dataset, model, and scaling method.
TABLE XX: Memory Usage (kB) per Dataset and Scaling Method
Dataset NO MM MA ZSN PS VAST MC RS QT DS TT LS HT
Breast Cancer
Wisconsin
Diagnostic
0.1875 176.8210 175.7480 176.2767 229.9746 229.1289 228.2461 384.9801 267.6338 229.8695 232.2894 230.8722 232.3309
Dry Bean 1704.2923 2388.0706 2448.4049 2599.3600 2404.5947 2405.1895 2404.3014 2388.2383 2520.3968 2406.4144 2406.9097 2405.5592 2408.1296
Glass
Identification
0.1875 23.2331 23.2064 26.1930 36.6759 36.0576 34.7992 51.7995 34.2393 36.1146 38.0356 37.9139 40.7885
Heart Disease 33.6901 49.0889 67.2025 71.7708 62.5277 62.8802 62.7256 121.9645 74.4512 63.1191 64.0603 64.0238 65.3331
Iris 0.1875 8.6296 8.7682 10.5488 16.9004 16.4834 15.6758 20.2601 22.0664 16.0218 18.4840 18.5913 20.3986
Letter
Recognition
0.1875 2567.9333 3566.3210 3787.2116 3629.7077 3629.3434 3628.4954 2568.3698 3318.4468 3628.8434 3632.1798 3631.9978 3742.5224
Magic Gamma
Telescope
0.1875 1552.4850 1552.0501 1552.2611 2198.9515 2199.2367 2198.2923 1552.7106 1792.9922 2198.5312 2201.3631 2201.1431 2306.3952
Rice Cammeo
And Osmancik
211.2500 297.4365 358.2734 378.0465 304.2510 304.3102 303.1006 297.6322 359.7692 304.2708 305.9161 305.7961 328.2706
Wine 21.0007 31.2210 40.3809 43.8314 45.0397 44.8984 44.5498 73.9326 46.5228 45.7445 46.4517 47.1243 47.9475
Abalone 0.1875 294.7493 294.4055 294.5703 355.6029 354.9523 353.4513 295.0916 360.0691 353.7972 356.8120 356.2417 380.5150
Air Quality 880.1009 1233.9790 1294.5547 1373.1566 1247.1017 1246.0932 1245.9476 1234.2925 1335.2915 1246.6934 1247.9374 1247.9843 1248.8872
Appliances
Energy
Prediction
4164.2812 5832.9988 5894.4865 6261.5162 5867.4751 5867.0134 5865.2385 5832.9714 6051.0226 5868.1863 5868.2539 5868.5447 5970.7936
Concrete
Compressive
Strength
65.6575 94.2771 137.7692 144.9830 104.1352 103.5251 102.7008 94.4553 181.1929 103.1220 105.1453 104.8095 112.0503
Forest Fires 43.2770 62.2524 87.1918 92.4231 72.8778 72.3182 71.9400 157.8339 105.3538 72.2997 72.9565 72.5387 74.4466
Real Estate
Valuation
22.2614 32.7211 43.1918 46.3411 38.4530 38.1743 36.8636 79.0026 67.4926 38.1566 40.1310 39.8900 43.6567
Wine Quality 0.1875 624.8189 624.3658 624.5845 833.4797 833.3540 831.6786 625.0661 744.7711 832.4357 834.7470 835.6562 871.7646
TABLE XX: Memory Usage (kB) per Dataset and Scaling Method

### v2

@@
 $X_{\text{norm}}=\frac{X-\mu}{\sigma},$ (3)
-where $X$ is the original feature value, $\mu$ is the mean of the feature, $\sigma$ is the standard deviation [^1] For transparency, we clarify that the population standard deviation was used throughout this study.
-of the feature, and $X_{\text{norm}}$ is the scaled value.
+where $X$ is the original feature value, $\mu$ is the mean of the feature, $\sigma$ is the standard deviation of the feature, and $X_{\text{norm}}$ is the scaled value.
 III-C4 Variable Stability Scaling (VAST)
@@
 This section presents the empirical results of our extensive experiments.
-We selected five models and three scaling techniques, along with a baseline without scaling, as this subset already allows us to draw meaningful conclusions and observe differences between models.
+We selected five representative models and three scaling techniques, along with a baseline without scaling, as this subset already allows us to draw meaningful conclusions and observe distinctions among models.
 The complete results are available in the Appendix A.
@@
 SVR 0.6076 0.5453 0.5111 0.5142
-As anticipated, a key finding from our classification experiments is the differential impact of feature scaling on model performance.
+As anticipated, one of the key findings from our classification experiments is the differential impact of feature scaling on model performance.
 Table III shows the accuracy results.
-Ensemble methods, including Random Forest and the gradient boosting family (LightGBM, CatBoost, XGBoost), proved robust, delivering consistently high validation metrics irrespective of the preprocessing strategy.
+Ensemble methods, including Random Forest and the gradient boosting family (LightGBM, CatBoost, XGBoost), demonstrated strong robustness by consistently achieving high validation performance irrespective of the preprocessing strategy or dataset.
 This inherent robustness offers a significant practical advantage, particularly in resource-constrained environments, since omitting the scaling step eliminates the associated memory and computational overhead.
 The Naive Bayes model showed similar scaling resistance, although its overall accuracy was not competitive with these top-tier ensembles.
-In stark contrast, the performance of Logistic Regression (LR), Support Vector Machines (SVM), K-Nearest Neighbor (KNN), TabNet and Multi-Layer Perceptrons (MLP) was highly dependent on the choice of scaler, revealing their pronounced sensitivity to data preprocessing through significant fluctuations in performance.
+In stark contrast, the performance of Logistic Regression (LR), Support Vector Machines (SVM), K-Nearest Neighbor (KNN), TabNet, and Multi-Layer Perceptrons (MLP) was highly dependent on the choice of scaler, revealing their pronounced sensitivity to data preprocessing through significant fluctuations in performance.
 A similar pattern of scaling sensitivity was observed in regression tasks when employing the regression counterparts of these classification models, as shows Tables IV, V and VI.
@@
 This practical consideration, combined with their predictive power, underscores their utility in a wide range of applications.
-IV-B Train and Inference time Impact
+IV-B Impact on Training and Inference Times
 TABLE VII: Time to Train (s) in classification models by dataset, model, and scaling method, showing only KNN, LGBM, MLP, RF, and SVM under the None, MA, ZSN, and RS scaling methods.
@@
 SVR 0.1501 0.1527 0.1419 0.1422
-The application of different scaling techniques had a variable impact on inference times in all models.
+The application of different scaling techniques had a variable impact on inference times across the evaluated models.
+Notably, Classification and Regression Trees (CART) exhibited exceptionally robust behavior, remaining unaffected across all scaling methods and datasets.
+While certain Machine Learning algorithms — such as K-Nearest Neighbors (KNN), Random Forest, Support Vector Machine (SVM), and Support Vector Regressor (SVR) — showed more evident sensitivity to the choice of scaling technique, this was not the norm.
+For the majority of models, the preprocessing step introduced only a small, non-uniform computational overhead — as illustrated in Tables VIII and X — which may become more significant in the context of large-scale datasets or time-sensitive applications requiring real-time inference.
 For training time, the effects of scaling largely mirrored those seen in the validation accuracy results, as shown in Tables VII and IX.
 Certain models, notably tree-based ensembles, did not derive a significant speed benefit from feature scaling, while others were more sensitive.
-The preprocessing step for inference generally introduced a small, non-uniform computational cost, which could become more pronounced for large datasets or in applications requiring real-time inference.
 IV-C Results of Memory Usage (kB)
-The analysis showed a clear distinction in memory consumption between scaling techniques.
+The analysis revealed a clear distinction in memory consumption between scaling techniques.
 As expected, applying no scaling resulted in zero additional memory consumption, only the consumption to load the data.
-Among the actual scaling methods, the RobustScaler, StandardScaler,Tanh Transformer, and Hyperbolic Tangent were found to be the most memory intensive.
-In contrast, the MaxAbsScaler, MinMaxScaler, and Decimal Scaler consistently registered the lowest memory usage as shows in Table XI.
+Among the actual scaling methods, the RobustScaler, StandardScaler, Tanh Transformer, and Hyperbolic Tangent were found to be the most memory-intensive.
+In contrast, the MaxAbsScaler, MinMaxScaler, and Decimal Scaler consistently registered the lowest memory usage as shown in Table XI.
 TABLE XI: Memory Usage (kB) per Dataset and Scaling Method
@@
 While this study provides a broad empirical analysis of feature scaling across various models and datasets, certain limitations should be acknowledged, which also open avenues for future research.
-  • Hyperparameter Optimization: The Machine Learning models in this study were evaluated using their default hyperparameters, as stated in the methodology.
-A comprehensive hyperparameter tuning process for each model-scaler-dataset combination was beyond the current scope but could potentially reveal different optimal pairings or further enhance model performance.
+  • Hyperparameter Optimization: The Machine Learning models analyzed in this study were evaluated using their default hyperparameters, as outlined in the methodology.
+A comprehensive hyperparameter tuning process for each model–scaler–dataset combination was beyond the current scope; however, such optimization could potentially uncover different optimal pairings or further improve model performance.
   • Scope and Diversity of Datasets: Although 16 datasets were used for both classification and regression tasks, the findings could be further enhanced by incorporating an even wider array of datasets, particularly those with very high dimensionality, different types of underlying data distributions, or from more specialized domains.
   • Evaluation Metrics for Classification: The primary metric for classification tasks was accuracy.
-Although acknowledged as potentially misleading for imbalanced datasets, future work could incorporate a broader suite of metrics, such as F1 score, precision-recall AUC, or balanced precision, to provide a more nuanced understanding of performance, especially on datasets with skewed class distributions.
-  • Dataset Size and Synthetic Data: For some of the smaller datasets utilized, the exploration of techniques like synthetic data generation or data augmentation was not performed.
-Such methods could potentially improve the robustness and performance of certain models, particularly data-hungry algorithms, and would be an interesting direction for further study.
+Although acknowledged as potentially misleading for imbalanced datasets, future work could incorporate a broader suite of metrics, such as F1 score, precision, recall AUC, or balanced precision, to provide a more nuanced understanding of performance, especially on datasets with skewed class distributions.
+  • Dataset Size and Synthetic Data: For some of the smaller datasets utilized, the exploration of techniques such as synthetic data generation or data augmentation was not performed.
+Such methods could potentially improve the robustness and performance of certain models, representing a promising direction for further research.
   • Focus on Default Algorithm Implementations: The study relied on standard implementations of algorithms mainly from well-known libraries.
@@
 VI Conclusion
-This comprehensive empirical study investigated the impact of 12 feature scaling techniques on 13 Machine Learning algorithms in 16 classification and regression datasets.
+This comprehensive empirical study investigated the impact of 12 feature scaling techniques on 14 Machine Learning algorithms in 16 classification and regression datasets.
 Key findings reaffirmed the robustness of ensemble methods (e.g., Random Forest, gradient boosting family), which, along with Naive Bayes, largely maintained high performance irrespective of scaling.
 This offers efficiency gains by potentially avoiding preprocessing overhead.
-In stark contrast, models such as logistic regression, SVMs, MLPs, K-Nearest Neighbor and TabNet demonstrated high sensitivity, with their performance critically dependent on scaler choice - a pattern consistent across both task types.
+In stark contrast, models such as Logistic Regression, SVMs, MLPs, K-Nearest Neighbor, and TabNet demonstrated high sensitivity, with their performance critically dependent on scaler choice — a pattern consistent across both task types.
 Computational analysis also indicated that scaling choices can influence training/inference times and memory usage, with certain scalers being notably more resource-intensive.
-This study contributes with one of the first systematic evaluations of such an extensive array of models, some less common models like TabNet, and scaling techniques—including transformations, e.g Tanh (TT) and Hyperbolic Tangent (HT), which are less commonly benchmarked as general-purpose scalers, all within a unified Python framework.
+This study contributes with one of the first systematic evaluations of such an extensive array of models, some less common models like TabNet, and scaling techniques — including transformations, e.g Tanh (TT) and Hyperbolic Tangent (HT), which are less commonly benchmarked as general-purpose scalers —, all within a unified Python framework.
 This is particularly relevant given that feature scaling is often applied in the literature without clear rationale, sometimes incorrectly before data splitting, which can lead to data leakage, or without verifying algorithm-specific benefits.